### v1

[…]
credit assignment, critical horizon, deep reasoning, process supervision, signal decay, inspection design
1 Introduction
When a semiconductor wafer fails final testing, which of the hundred preceding fabrication steps caused the defect? When a customer churns after a multi-touchpoint journey, which interaction was responsible? When a multi-tier supply chain delivers a defective product, which supplier introduced the flaw? These are instances of the credit assignment problem: attributing terminal outcomes to intermediate stages in deep sequential processes.
Classical inspection theory (Lindsay and Bishop 1964, Raz 1986, Tapiero 1987) addresses this challenge through intermediate checkpoints, optimizing their placement to balance inspection cost against defect-escape risk. But intermediate inspection is not always an option: a proof step that appears irregular may be valid in context; a component may only be testable once assembled; a customer’s reaction to one touchpoint depends on what follows. In such cases, validity is a property of the whole sequence, and terminal inspection is not merely cheaper but necessary. Moreover, even when terminal inspection is the only option, the classical literature implicitly assumes that defect signals persist through the production chain: with enough terminal observations, any stage’s quality can eventually be inferred. We show this assumption fails for deep processes, where signals decay exponentially with depth.
The challenge of attributing outcomes to early stages under terminal-only observation has gained particular urgency in AI, where large language models generate extended reasoning chains but only the final answer can be verified. Empirically, value networks trained on outcome rewards “collapse”: they converge to near-constant predictions for early steps, as if outcomes carry no information about the chain’s beginning (Kazemnejad et al. 2024). Process reward models that provide step-level supervision dramatically improve performance (Lightman et al. 2024), but require expensive human annotation. When is such annotation necessary, and when does outcome supervision suffice?
This paper answers these questions by establishing information-theoretic foundations for credit assignment. Using Strong Data Processing Inequalities, we prove that signals connecting early steps to terminal outcomes decay exponentially with depth. The analysis reveals a phase transition: there exists a critical horizon $H_{\mathrm{crit}}$ such that steps within this distance of the outcome can be evaluated from endpoint data, while steps beyond cannot, regardless of sample size, algorithm, or computational budget. This is not an algorithmic limitation but an information-theoretic impossibility. The critical horizon scales logarithmically with sample size and signal strength but inversely with a per-stage contraction rate that captures how much each processing step blurs the distinction between good and bad earlier steps. In AI reasoning systems, sampling temperature controls this blurring: higher temperature produces more diverse continuations, making it harder to trace outcomes back to early steps. In manufacturing, the analogous driver is the overlap among downstream quality-grade distributions: variability-reducing steps (polishing, annealing, rework), pooling operations (lot mixing, kitting, assembly), and coarse metrology bins all cause products of different upstream grades to produce similar distributions of grades at the next stage, fading the signal about the original grade. The critical horizon formula thus determines whether intermediate feedback is a cost-saving measure or a feasibility requirement.
A key insight underlying our analysis is that credit assignment targets abstract states, not raw observations. In manufacturing, the observable state at production stage $t$ consists of detailed sensor readings (temperatures, pressures, vibration spectra), but credit assignment asks whether the stage produced output that was within specification, marginal, or defective. Multiple sensor patterns correspond to the same quality grade. In AI reasoning, the observable state is the token sequence generated after $t$ reasoning steps, but credit assignment asks whether the logical content is valid, not which specific words were used. The phrases “Therefore x=5x=5” and “Thus we conclude $x$ equals 55” use different tokens but express the same mathematical content. In both settings, the space of abstract states (quality grades, semantic content) is smaller than the space of raw observations (sensor readings, token sequences).
This many-to-one mapping from observations to abstract states is the mechanism behind the per-stage contraction introduced above: when multiple observable states collapse to the same abstract state, transitions from distinct starting points lead to overlapping distributions over next states, and statistical distinguishability shrinks. This occurs even when raw dynamics are deterministic (as with token generation in language models), because the induced dynamics on the abstract space contract. The implication for learning systems is immediate: when signal decay is severe, outcome distributions conditioned on different early states become nearly identical, and training data contains almost no information about which early states are better. The critic collapse phenomenon observed in AI systems (Kazemnejad et al. 2024) is thus statistically optimal behavior given the available information, not a bug to be fixed but a symptom of a fundamental limit.
A natural response to signal decay is to generate more observations: multiple parallel trajectories from the same starting point, as in replicated testing or width-based methods like GRPO (DeepSeek-AI 2024). We show this strategy provides only limited relief: correlation among rollouts caps the effective width at $1/\rho$, where $\rho$ is the pairwise correlation between rollout outcomes, and adding more rollouts yields diminishing returns.
For steps beyond the critical horizon, process supervision becomes necessary, raising a new question: how should step-level feedback be aggregated? Sequential validity requires all steps to be correct, a multiplicative rather than additive criterion. The mismatch is stark: a 100-step process with 99% per-step accuracy has an expected correctness score of 99, yet the probability that all 100 steps are correct is only $0.99^{100}\approx 0.37$. Systems that optimize additive objectives learn to produce chains that are “mostly correct” but nonetheless invalid. We analyze this objective mismatch and propose curriculum strategies that transition from learnable additive objectives to correct multiplicative ones as policy quality improves.
Once the need for process supervision is established, a natural question arises: where should intermediate checkpoints be placed? Each checkpoint resets the signal decay clock, so a step immediately before a checkpoint is easy to evaluate, while a step far from any checkpoint is hard. The hardest step to evaluate under any schedule is the one farthest from its next checkpoint, and the design goal is to make this worst case as mild as possible. We derive inspection schedules that are minimax-optimal in this sense: they minimize the worst-case sample complexity across all steps. When all stages attenuate signal at the same rate, uniform checkpoint spacing is minimax-optimal. When different stages attenuate at different rates, a greedy algorithm that places each successive checkpoint as far as the information budget allows yields optimal non-uniform schedules, automatically concentrating inspections where signal loss is fastest. In either case, achieving polynomial sample complexity requires inspection density that scales nearly linearly with process depth: deeper processes demand proportionally more intermediate verification.
We develop a detailed application to large language model reasoning (Section 14 in the Electronic Companion), where the state abstraction maps token sequences to semantic content and sampling temperature controls the contraction coefficient. We validate the theoretical predictions through numerical experiments on both synthetic Markov chains and LLM reasoning chains on GSM8K (a grade-school math benchmark).
Summary of contributions.
We establish four main results that transform the intuition “credit assignment in long processes is hard” into precise, actionable principles.
  • Signal Decay Bound (Section 3): Information about any given step decays exponentially when propagating to the terminal outcome. The resulting sample complexity bounds are tight, establishing that exponential scaling is fundamental. This yields a closed-form critical horizon formula.
  • Width Limits (Section 4): Parallel rollouts reduce estimator variance, but correlation caps effective width at $1/\rho$, extending the critical horizon only logarithmically.
  • Objective Mismatch (Section 5): Additive reward aggregation optimizes the wrong quantity when sequential validity requires all steps to be correct. We propose curriculum strategies that interpolate between learnable and correct objectives.
  • Optimal Inspection Design (Section 6): Uniform checkpoint spacing is minimax-optimal under homogeneous contraction; a greedy algorithm yields optimal non-uniform schedules under heterogeneous contraction. Polynomial sample complexity requires inspection density scaling as $\Omega(H/\log H)$.
The framework yields concrete guidance for both AI and operations practitioners. For AI systems, it explains why process reward models improve performance (they provide signal beyond the critical horizon), why methods like VinePPO (Kazemnejad et al. 2024) succeed (they reduce effective depth below $H_{\mathrm{crit}}$), and why width-based methods face diminishing returns (correlation caps effective width). For operations managers, it provides a direct diagnostic. Consider a 50-stage semiconductor process. With contraction rate $\eta=0.85$, signal strength $\Delta^{2}=0.2$, and 10,000 test wafers, $H_{\mathrm{crit}}\approx 47$: at least one intermediate checkpoint is required. With $\eta=0.90$, the critical horizon extends to 72 steps, and terminal testing alone suffices. (In practice, $\eta$ is auditable: estimate adjacent-stage grade transition matrices from production data and compute the overlap-based bounds in Section 3.2.)
The unifying insight is that credit assignment difficulty is a structural barrier, not an algorithmic one. The critical horizon formula answers the diagnostic question (when is intermediate feedback necessary?); the optimal inspection schedules answer the prescriptive question (where to place checkpoints?); and the objective mismatch analysis answers the aggregation question (how to combine step-level feedback?). Together, these results provide a principled foundation for supervision system design in both operations and AI.
1.1 Literature Review
The information-theoretic structure underlying credit assignment connects five research streams that have developed largely independently: inspection design in operations research, credit assignment in AI reasoning, reinforcement learning theory, information-theoretic analysis of sequential processes, and the emerging application of OR tools to AI system design.
Inspection design in operations research.
The optimal placement of inspection points in sequential production is a foundational topic in operations research. Lindsay and Bishop (1964) formulated inspection allocation as a serial dynamic program, minimizing total cost of inspection plus penalties for escaping defects. Raz (1986) surveyed extensions incorporating multiple inspection stations and variable costs; Tapiero (1987) introduced imperfect inspection and learning effects. Recent work incorporates adaptive inspection (Tan et al. 2024), predictive maintenance (Compare et al. 2020), and machine-learning-based prognosis (Diez-Olivan et al. 2019). Ghuge et al. (2025) study sequential testing of system components under budget constraints, providing the first constant-factor approximation for nonadaptive score classification. Duenyas et al. (2024) analyze supply chains where components cannot be individually tested, developing coordination mechanisms when stage-level inspection is infeasible, precisely the terminal-only observation regime our theory addresses.
This literature optimizes checkpoint placement to minimize cost under the assumption that defect signals persist through the production chain. Our work identifies a structural limit: in deep processes, signal decay makes observability, not cost, the binding constraint. The critical horizon formula quantifies when this regime applies.
Credit assignment in AI reasoning.
The credit assignment problem was recognized by Minsky (1961) as fundamental to learning systems, yet has resisted systematic theoretical treatment. Our framework provides an information-theoretic characterization of when outcome supervision suffices versus when process supervision becomes necessary, grounded in the signal decay structure of deep sequential processes.
Recent work explores process supervision as a remedy: Lightman et al. (2024) show that step-level labels substantially improve mathematical reasoning; Uesato et al. (2022) compare outcome and process supervision for mathematical reasoning; Wang et al. (2024) develop automated step-labeling methods. These studies demonstrate that process supervision helps but do not characterize when it becomes necessary. Our Signal Decay Bound fills this gap.
Empirically, value networks trained on outcome data exhibit “critic collapse”, converging to near-constant predictions on long chains (Kazemnejad et al. 2024). Width-based methods such as GRPO (DeepSeek-AI 2024) address this by replacing learned critics with Monte Carlo estimation. Our framework explains both phenomena: critic collapse reflects the absence of signal beyond the critical horizon, and width-based methods face diminishing returns because correlation caps effective width.
Reinforcement learning theory.
Standard reinforcement learning theory establishes polynomial sample complexity under regularity conditions such as bounded concentrability (Kakade 2003, Azar et al. 2013). Recent work by Jia et al. (2025) shows that outcome and process supervision are polynomially equivalent when concentrability is bounded. Our setting falls outside this regime: in deep sequential processes where each step may fail, concentrability grows exponentially in horizon, and the polynomial equivalence breaks down. This is precisely why the critical horizon matters. Sample complexity in related RL settings has received attention in the OR community: Hu et al. (2025) establish fast convergence rates for offline RL, Cheung et al. (2023) develop algorithms for nonstationary environments, and Lykouris et al. (2025) address exploration under adversarial corruption. These works assume bounded concentrability or benign nonstationarity; the deep sequential regime we study, where signal decays exponentially, falls outside their scope.
A second point of departure from standard reinforcement learning concerns objective structure. Standard algorithms aggregate rewards additively, but sequential validity requires all steps to be correct, inducing a multiplicative objective. This connects to risk-sensitive MDPs (Howard and Matheson 1972), multiplicative-reward models (Baier et al. 2025), and minimax-optimal risk-sensitive RL (Wang et al. 2023). We propose a curriculum strategy that interpolates between the learnable additive objective and the correct multiplicative one.
Information-theoretic foundations.
Filter stability results (Del Moral and Guionnet 2001) establish that dependence on initial conditions decays asymptotically in Markov chains, but do not yield the non-asymptotic, closed-form rates that credit assignment requires: a practitioner needs to know how many samples suffice for a process of a given depth, not merely that the problem becomes harder in the limit. Our analysis provides such rates by combining two classical tools. Strong Data Processing Inequalities (Polyanskiy and Wu 2017, Raginsky 2016) quantify per-step information loss, establishing that $\chi^{2}$-divergence contracts by a factor $\eta<1$ at each transition. Le Cam’s method (Le Cam 1973) then translates the resulting divergence bounds into sample complexity lower bounds, yielding the critical horizon formula. Recent extensions to bandits under communication constraints (Pensia et al. 2024) and interactive learning (Chen et al. 2024) address related but distinct settings.
Operations research tools for AI system design.
A growing body of work applies OR methodology directly to AI systems. Snell et al. (2025) frame inference-time compute allocation as an optimization problem, adaptively distributing verification effort across prompts of varying difficulty, an allocation problem that parallels inspection placement at the problem-structure level, though without the information-theoretic signal decay analysis we develop here. Besta et al. (2025) formulate LLM reasoning explicitly within the MDP framework, treating chain-of-thought generation as sequential decision-making. Zhang et al. (2024) apply Monte Carlo Tree Search to guide multi-step reasoning with step-level evaluation, and Tang et al. (2024) model iterative code repair as a multi-armed bandit. Pieroth et al. (2025) use deep RL to compute equilibrium strategies in multistage auctions, and Elmachtoub and Grigas (2022) establish the “predict, then optimize” framework integrating prediction with downstream decision-making, both demonstrating OR-native methodology enriching AI system design. Recent surveys of INFORMS Fellows identify bringing OR rigor to AI challenges as a priority research frontier (Wiberg et al. 2025). Our work contributes to this agenda by showing that Strong Data Processing Inequalities and minimax inspection design, classical OR tools, yield precise, actionable answers to questions the AI community has addressed primarily through empirical heuristics.
1.2 Paper Organization
The paper develops a logical progression from information-theoretic limits to optimal inspection design. Section 2 establishes the formal framework: the MDP model, state abstraction, supervision regimes, and the contraction property that drives signal decay. Section 3 proves the Signal Decay Bound, showing that information about step $t$ decays as $\eta^{H-t}$, and derives the critical horizon formula that determines which steps can be evaluated from terminal data.
Two extensions follow. Section 4 analyzes width amplification, showing that correlation among parallel rollouts caps effective width and extends the critical horizon only logarithmically. Section 5 addresses an orthogonal failure mode: standard additive objectives optimize the wrong quantity for sequential validity, and proposes curriculum strategies that interpolate between learnable and correct objectives.
Section 6 synthesizes these results into optimal inspection schedules: uniform spacing under homogeneous contraction and a greedy information-distance algorithm under heterogeneous contraction. Section 7 validates the theoretical predictions through synthetic Markov chains and LLM reasoning chains on GSM8K, covering all four main results. The Electronic Companion develops a detailed application to language model reasoning (Section 14).
2 Problem Formulation
We formalize the credit assignment problem in sequential processes, beginning with the operational setting before developing the mathematical framework.
2.1 Multi-Stage Processes with Terminal Inspection
Consider a process where items pass through $H$ stages before final inspection. At each stage, the item’s state changes and defects may be introduced. The final inspection reveals whether the complete item meets specifications but does not identify which stage caused a failure.
As discussed in Section 1, terminal-only inspection arises either because intermediate inspection is costly (semiconductor metrology, supplier auditing) or infeasible (validity depends on context that emerges only downstream). The common structure across manufacturing, service operations, supply chains, and AI reasoning is that terminal outcomes are observable but the intermediate stage responsible for success or failure is not. Without localization, process improvement efforts may target the wrong stages.
Our framework applies when (i) the process evolves through discrete stages with Markovian dynamics on an abstract state space, (ii) terminal inspection reveals aggregate quality, and (iii) intermediate inspection, if feasible, is costly. In AI, outcome supervision refers to training on final output labels, while process supervision provides step-level feedback. We adopt this terminology throughout, noting that the underlying structure is domain-general.
2.2 Mathematical Formalization
We model the sequential process as an episodic finite-horizon Markov decision process (MDP) with terminal reward. Let $\mathcal{S}$ denote a state space, $\mathcal{A}$ an action space, and $H$ the horizon (the number of decision stages).
Definition 2.1 (Finite-Horizon MDP with Terminal Reward)
A finite-horizon MDP with terminal reward consists of:
  • An initial state $S_{0}\in\mathcal{S}$, drawn from a distribution $\mu$ over $\mathcal{S}$ (or fixed at a given state $s_{0}$)
  • Transition dynamics, either deterministic $s_{t+1}=f(s_{t},a_{t})$ or stochastic $s_{t+1}\sim P(\cdot\mid s_{t},a_{t})$
  • A policy $\pi=(\pi_{0},\ldots,\pi_{H-1})$ where each $\pi_{t}:\mathcal{S}\to\Delta(\mathcal{A})$ maps states to distributions over actions
  • A terminal evaluation $R=g(S_{H})\in\{0,1\}$ indicating success or failure
  • Step validity indicators $r_{t}=h_{t}(S_{t},A_{t},S_{t+1})\in\{0,1\}$, treated as exogenous labels available only under process or partial supervision; the functions $h_{t}$ are problem-dependent and need not be known to the learner
A trajectory is the complete sequence $\tau=(S_{0},A_{0},S_{1},\ldots,A_{H-1},S_{H})$. We use uppercase letters for random variables; realized samples are denoted $\tau^{(i)}=(s_{0}^{(i)},a_{0}^{(i)},\ldots,s_{H}^{(i)})$.
To ground this abstraction: in manufacturing, the state $S_{t}$ is the workpiece condition at stage $t$, an action $A_{t}$ is the process setting (temperature, pressure, feed rate), and the transition $f$ maps current condition and setting to the next-stage condition. The terminal evaluation $R$ indicates whether the finished product passes final inspection. In AI reasoning, $S_{t}$ is the partial solution generated so far, $A_{t}$ is the next reasoning step, and $R$ indicates whether the final answer is correct.
Remark 2.2 (Deterministic Dynamics with Stochastic Trajectories)
In many sequential processes, the transition function $f$ is deterministic while trajectories are stochastic because actions are sampled from a policy distribution. This distinction matters: information loss arises not from noisy dynamics but from the many-to-one structure of state abstraction (Section 3). In manufacturing, this occurs even under fixed recipes: units sharing the same quality grade may differ in latent characteristics (material lot, tool drift, defect mode), so the abstraction from detailed workpiece state to grade is many-to-one, and the induced grade-to-grade kernels have overlapping rows and $\eta<1$.
The value function captures the probability of eventual success from any intermediate state. Intuitively, it answers the question: “If the process is currently at state $s$ after $t$ steps, what is the chance it will ultimately succeed?”
Definition 2.3 (Value Function)
The value function under policy $\pi$ from state $s$ at step $t$ is
$V^{\pi}_{t}(s):=\mathbb{P}_{\pi}(R=1\mid S_{t}=s)=\mathbb{E}_{\pi}[R\mid S_{t}=s],$ (1)
representing the probability of terminal success when starting from state $s$ at step $t$ and following policy $\pi$.
2.3 Supervision Models
The learner observes $n$ i.i.d. trajectories generated by a fixed behavior policy $\pi$. We define three supervision regimes, distinguished by what feedback accompanies each trajectory.
Definition 2.4 (Outcome Supervision)
Under outcome supervision, the learner observes $n$ independent trajectory-outcome pairs:
$\mathcal{D}_{\mathrm{out}}=\{(\tau^{(i)},R^{(i)})\}_{i=1}^{n}.$ (2)
The full trajectory $\tau^{(i)}$ including all intermediate states is observed, but no step-level correctness labels are provided.
The learner sees what happened at each step but not whether each step was correct.
Definition 2.5 (Process Supervision)
Under process supervision, the learner additionally observes step validity labels:
$\mathcal{D}_{\mathrm{proc}}=\{(\tau^{(i)},r_{0}^{(i)},\ldots,r_{H-1}^{(i)},R^{(i)})\}_{i=1}^{n},$ (3)
where $r_{t}^{(i)}\in\{0,1\}$ indicates whether step $t$ of trajectory $i$ is valid.
Process supervision provides the ground truth about each step. This is the gold standard for credit assignment but is expensive: a human expert must evaluate every step of every trajectory, or a formal verifier must check each logical inference.
Definition 2.6 (Partial Supervision)
Under $k$-step supervision, the learner observes step labels at intervals of $k$ steps:
$\mathcal{D}_{k}=\{(\tau^{(i)},r_{k}^{(i)},r_{2k}^{(i)},\ldots,r_{\lfloor(H-1)/k\rfloor\cdot k}^{(i)},R^{(i)})\}_{i=1}^{n},$ (4)
where the labeled steps are $\{t\in\{1,\ldots,H-1\}:t\equiv 0\pmod{k}\}$. This interpolates between near-process supervision ($k=1$, all steps except $r_{0}$ labeled) and outcome supervision ($k\geq H$, no step labels). Omitting $r_{0}$ is without loss for our results; it ensures that coarse supervision provides no step-level information.
Under all three regimes, the learner observes the complete state sequence $(s_{0},\ldots,s_{H})$. The difficulty is not hidden states but missing correctness labels: the model’s output is fully readable, but determining whether each step is valid may require human experts or formal verification.
2.4 Information-Theoretic Preliminaries
Our lower bounds rely on tracking how information degrades through sequential processing. The core question is: if two hypotheses (say, “step $t$ was good” versus “step $t$ was bad”) lead to different distributions over trajectories, how many samples suffice to distinguish them?
The answer depends on how different the resulting outcome distributions are: nearly identical distributions require many samples, while well-separated distributions require few. We formalize this using $\chi^{2}$-divergence.
Definition 2.7 ($\chi^{2}$-Divergence)
For distributions $P$ and $Q$ with $P$ absolutely continuous with respect to $Q$:
$\chi^{2}(P\|Q):=\int\frac{(dP-dQ)^{2}}{dQ}=\mathbb{E}_{Q}\left[\left(\frac{dP}{dQ}-1\right)^{2}\right].$ (5)
We use $\chi^{2}$-divergence rather than KL divergence for two reasons. First, it connects directly to hypothesis testing: by Le Cam’s method (Le Cam 1973), distinguishing $P$ from $Q$ with constant probability requires $n=\Omega(1/\chi^{2}(P\|Q))$ samples. Second, it tensorizes cleanly across independent samples: $\chi^{2}(P^{\otimes n}\|Q^{\otimes n})=(1+\chi^{2}(P\|Q))^{n}-1$ (see, e.g., Polyanskiy and Wu 2017). This multiplicative structure translates divergence bounds directly into sample complexity bounds.
The central tool in our analysis is the Strong Data Processing Inequality, which captures a fundamental phenomenon: processing information can only destroy it, never create it. If two distributions are hard to distinguish, passing them through a noisy channel makes them even harder to distinguish.
Definition 2.8 ($\chi^{2}$-Contraction Coefficient)
For a Markov kernel $K:\mathcal{X}\to\Delta(\mathcal{Y})$, the $\chi^{2}$-contraction coefficient is
$\eta_{\chi^{2}}(K):=\sup_{\begin{subarray}{c}P,Q:P\ll Q\\ 0<\chi^{2}(P\|Q)<\infty\end{subarray}}\frac{\chi^{2}(PK\|QK)}{\chi^{2}(P\|Q)},$ (6)
where $PK(y)=\int K(y\mid x)\,dP(x)$ is the pushforward measure.
The contraction coefficient satisfies $\eta_{\chi^{2}}(K)\in[0,1]$, measuring what fraction of distinguishability survives the channel: $\eta=0$ destroys all information; $\eta=1$ is lossless. Throughout, smaller $\eta$ means stronger contraction (faster information loss).
Lemma 2.9 ($\chi^{2}$ Strong Data Processing Inequality (see Raginsky 2016, Polyanskiy and Wu 2017))
Let $K:\mathcal{Y}\to\Delta(\mathcal{Z})$ be a Markov kernel. For any distributions $P_{Y}$ and $Q_{Y}$ on $\mathcal{Y}$ with $P_{Y}\ll Q_{Y}$ and $\chi^{2}(P_{Y}\|Q_{Y})<\infty$, define $P_{Z}=P_{Y}K$ and $Q_{Z}=Q_{Y}K$. Then
$\chi^{2}(P_{Z}\|Q_{Z})\leq\eta_{\chi^{2}}(K)\cdot\chi^{2}(P_{Y}\|Q_{Y}).$ (7)
For credit assignment: if $P_{Y}$ and $Q_{Y}$ represent state distributions under “good step” versus “bad step,” and $K$ represents subsequent reasoning, then outcome distributions are at most $\eta$ times as distinguishable as the original states. When we iterate through multiple steps, information loss compounds.
Corollary 2.10 (Divergence Decay in Markov Chains (cf. Polyanskiy and Wu 2017))
Consider a Markov chain $(S_{0},S_{1},\ldots,S_{H})$ with transition kernels $K_{0},K_{1},\ldots,K_{H-1}$. For any two initial distributions $P_{S_{0}}$ and $Q_{S_{0}}$ with $P_{S_{0}}\ll Q_{S_{0}}$ and $\chi^{2}(P_{S_{0}}\|Q_{S_{0}})<\infty$:
$\chi^{2}(P_{S_{H}}\|Q_{S_{H}})\leq\left(\prod_{t=0}^{H-1}\eta_{\chi^{2}}(K_{t})\right)\chi^{2}(P_{S_{0}}\|Q_{S_{0}}).$ (8)
In particular, if $\eta_{\chi^{2}}(K_{t})\leq\eta<1$ for all $t$, then
$\chi^{2}(P_{S_{H}}\|Q_{S_{H}})\leq\eta^{H}\cdot\chi^{2}(P_{S_{0}}\|Q_{S_{0}}).$ (9)
3 The Signal Decay Bound: Why Early Steps Are Hardest
We now develop our first main result: the Signal Decay Bound. In settings ranging from manufacturing quality control to service analytics, a decision-maker observes process outcomes and must infer which intermediate stages were responsible for success or failure. The central question is: how much signal about an early step survives propagation through the remaining stages? We prove that this signal decays exponentially with distance from the step to the outcome, making credit assignment to early steps fundamentally harder than to late ones.
The section proceeds in three steps. Section 3.1 introduces the state abstraction that maps observable states to a coarser space on which the dynamics are Markov. Section 3.2 shows that the many-to-one structure of this abstraction induces contraction in the abstract transition kernels, quantifies the contraction coefficient, and states the standing assumption for subsequent results. Section 3.3 then proves the Signal Decay Bound, derives the critical horizon, and establishes tightness. Proofs of all results in this section appear in Appendix 10.
3.1 State Abstraction and Abstract Dynamics
Credit assignment targets abstract content, not raw observations. In many sequential processes, the full observable state contains details (surface representation, sensor noise, formatting choices) that are irrelevant to the outcome. Multiple observable states may map to the same quality grade or semantic content. This many-to-one structure is the source of information loss, and formalizing it is the first step toward quantifying signal decay.
Definition 3.1 (Markov State Abstraction)
A Markov state abstraction is a mapping $\phi:\mathcal{S}\to\mathcal{Z}$ from observable states to an abstract state space $\mathcal{Z}$ satisfying the lumpability condition: for all $t$ and all $s,s^{\prime}\in\mathcal{S}$ with $\phi(s)=\phi(s^{\prime})=z$,
$\mathcal{L}(\phi(S_{t+1})\mid S_{t}=s)=\mathcal{L}(\phi(S_{t+1})\mid S_{t}=s^{\prime}).$ (10)
The abstract state at step $t$ is $Z_{t}=\phi(S_{t})$.
The lumpability condition, due to Kemeny and Snell (1960), ensures that the abstract process $(Z_{0},Z_{1},\ldots,Z_{H})$ is itself Markov: the distribution of the next abstract state depends only on the current abstract state, not on which observable state within the equivalence class was realized. When this condition holds, there exist well-defined abstract transition kernels $K_{t}:\mathcal{Z}\to\Delta(\mathcal{Z})$ with $K_{t}(z^{\prime}\mid z)=\mathbb{P}(\phi(S_{t+1})=z^{\prime}\mid\phi(S_{t})=z)$.
Lumpability is a modeling choice rather than a restrictive hypothesis: one defines $\phi$ so that the abstract dynamics are Markov. If an initial choice of $\phi$ fails the condition, one refines the partition by augmenting $\phi$ with additional context until the condition holds, at the cost of a larger abstract space. The interesting question is not whether lumpability holds (it can always be arranged) but how much the resulting abstract kernels $K_{t}$ mix: do different abstract states lead to overlapping or distinct distributions over successor states? A coarser abstraction lumps more observable states together, increasing overlap and accelerating information loss. A finer abstraction preserves more distinctions but may require a larger state space. Section 3.2 formalizes this tradeoff. Section 14 in the Electronic Companion develops a concrete instantiation for language model reasoning.
Example 3.2 (State Abstraction in Applications)
  • Manufacturing: The observable state $S_{t}$ consists of sensor readings at production stage $t$ (temperatures, pressures, vibration spectra). The abstract state $Z_{t}=\phi(S_{t})$ captures the quality grade: “within specification,” “marginal,” or “defective.” Multiple sensor patterns map to the same grade.
  • AI reasoning: The observable state $S_{t}$ is the token sequence generated after $t$ reasoning steps. The abstract state $Z_{t}=\phi(S_{t})$ captures semantic content: the mathematical claims established, the proof strategy adopted, or the intermediate values derived. Multiple token sequences expressing the same content map to the same abstract state.
The lumpability condition requires exact agreement of abstract transition probabilities within each equivalence class. In practice, this may hold only approximately: two sensor patterns classified as the same quality grade may lead to slightly different distributions over downstream grades. We return to this point in Section 3.3, after establishing the main bound, where we show that approximate lumpability incurs only a bounded, typically negligible penalty.
3.2 Why Abstract Dynamics Contract
The many-to-one structure of $\phi$ introduced in Section 3.1 causes information loss whenever distinct abstract states lead to overlapping next-state distributions. We now quantify this loss per step using the $\chi^{2}$-contraction coefficient $\eta_{\chi^{2}}(K)\in[0,1]$ (Definition 2.8), which measures what fraction of $\chi^{2}$-divergence survives when distributions pass through a kernel $K$. When different abstract states $z$ and $z^{\prime}$ produce overlapping distributions $K(\cdot\mid z)$ and $K(\cdot\mid z^{\prime})$ over next states, $\eta_{\chi^{2}}(K)<1$ and any statistical difference between two distributions over $Z_{t}$ shrinks after one transition.
To build intuition, suppose we want to determine whether the abstract chain was in state $z$ or state $z^{\prime}$ at step $t$ by observing the state at step $t+1$. If $K(\cdot\mid z)$ and $K(\cdot\mid z^{\prime})$ have disjoint support, observing $Z_{t+1}$ perfectly identifies $Z_{t}$, and no information is lost. But if both distributions assign positive probability to some state $y$, then observing $Z_{t+1}=y$ leaves residual uncertainty about whether $Z_{t}$ was $z$ or $z^{\prime}$. The greater the overlap between $K(\cdot\mid z)$ and $K(\cdot\mid z^{\prime})$, the more information is lost at each step.
Example 3.3 (Contraction Coefficient in Extreme Cases)
  • Deterministic dynamics ($\eta=1$): If each state $z$ transitions to a unique successor, the transition matrix is a permutation and $K(\cdot\mid z)$ and $K(\cdot\mid z^{\prime})$ have disjoint support for $z\neq z^{\prime}$. No information is lost: $\eta_{\chi^{2}}(K)=1$.
  • Uniform mixing ($\eta=0$): If $K(\cdot\mid z)=\nu$ for all $z$ (every state transitions to the same distribution $\nu$), then $Z_{t+1}$ is independent of $Z_{t}$ and all information about the past is destroyed in one step: $\eta_{\chi^{2}}(K)=0$.
  • Intermediate case: Consider a two-state chain with $K(1\mid 1)=1-p$, $K(2\mid 1)=p$, $K(1\mid 2)=p$, $K(2\mid 2)=1-p$. The overlap between rows is $2\min\{p,1-p\}$, and $\eta_{\chi^{2}}(K)=(1-2p)^{2}$. When $p=0$, the chain is deterministic and $\eta=1$; when $p=1/2$, both rows are identical and $\eta=0$.
The following two propositions formalize the relationship between row overlap and contraction, providing concrete sufficient conditions for $\eta<1$.
Proposition 3.4 (Contraction from Diversity)
Consider a transition kernel $K$ on an abstract state space with $m$ states. If $K$ spreads probability such that $K(z^{\prime}\mid z)\geq p_{\min}>0$ for all state pairs $(z,z^{\prime})$, then $\eta_{\chi^{2}}(K)\leq 1-m\cdot p_{\min}$.
Proposition 3.4 provides a sufficient but strong condition: requiring $K(z^{\prime}\mid z)\geq p_{\min}$ for all state pairs is often unrealistic. The following proposition gives a weaker condition based on the Dobrushin coefficient, requiring only that each pair of rows has some overlap.
Proposition 3.5 (Contraction from Row Overlap)
For a transition kernel $K$ on a finite abstract state space $\mathcal{Z}$, define the Dobrushin coefficient:
$\alpha(K):=\min_{z,z^{\prime}\in\mathcal{Z}}\sum_{y\in\mathcal{Z}}\min\{K(y\mid z),K(y\mid z^{\prime})\}.$
Then $\eta_{\chi^{2}}(K)\leq 1-\alpha(K)$.
The resulting bound on $\eta$ is looser than Proposition 3.4 but applies more broadly, since it requires only pairwise row overlap rather than full minorization. In practice, $\alpha(K)$ can be estimated by sampling continuations from different prefixes and measuring their distributional overlap. In manufacturing, the same bound is computed directly from shop-floor data: bin each stage’s output into grades via $\phi$, estimate $K$ from empirical adjacent-stage transition frequencies, and compute $\alpha(K)$ from the resulting matrix.
Example 3.6 (Manufacturing Grade Transitions)
Consider a three-grade abstraction $Z_{t}\in\{\mathrm{A},\mathrm{B},\mathrm{C}\}$ where $\mathrm{A}$ denotes “within specification,” $\mathrm{B}$ denotes “marginal/reworkable,” and $\mathrm{C}$ denotes “defective,” in a setting with rework and defect masking. A downstream step such as polishing followed by retest can partially repair marginal units while occasionally degrading good ones. A representative grade-to-grade kernel is
$K=\begin{pmatrix}0.85&0.14&0.01\\ 0.55&0.35&0.10\\ 0.20&0.30&0.50\end{pmatrix}$,
where rows index the incoming grade and columns the outgoing grade. The minimum pairwise row overlap is achieved by the $\mathrm{A}$ and $\mathrm{C}$ rows:
[…]
so $\alpha(K)=0.35$ and Proposition 3.5 gives $\eta_{\chi^{2}}(K)\leq 1-\alpha(K)=0.65$. Even when the underlying physical transformation is repeatable, coarse grading and partial defect masking produce strict per-stage contraction on the abstract state space.
The preceding examples and propositions show that contraction ($\eta<1$) arises whenever the rows of the abstract kernel have nontrivial overlap. The only abstract chains that avoid contraction entirely are those whose transition matrix is a permutation, so that every state has a unique successor. We now state this as a standing assumption.
{assumption}
[Contractive Dynamics] At each step $u$, the abstract transition kernel $K_{u}$ satisfies $\eta_{\chi^{2}}(K_{u})\leq\eta$ for some $\eta<1$.
Propositions 3.4 and 3.5 provide concrete upper bounds on $\eta$ in terms of the abstract kernel’s row overlap; Section 14 in the Electronic Companion shows how sampling temperature controls $\eta$ in language model reasoning.
Under Assumption 3.6, consider two hypotheses $H_{0}$ and $H_{1}$ that induce different distributions $P^{(0)}_{Z_{t}}$ and $P^{(1)}_{Z_{t}}$ over the abstract state at step $t$. Information about step $t$ must propagate through HtH-t transitions to reach the outcome:
$Z_{t}\xrightarrow{K_{t}}Z_{t+1}\xrightarrow{K_{t+1}}\cdots\xrightarrow{K_{H-1}}Z_{H}\to R.$
At each transition, the SDPI contracts the divergence by at most factor $\eta$, yielding:
$\text{At step }t:$ $\quad\chi^{2}(P^{(0)}_{Z_{t}}\|P^{(1)}_{Z_{t}})=\Delta^{2}$
$\text{At step }t+1:$ $\quad\chi^{2}(P^{(0)}_{Z_{t+1}}\|P^{(1)}_{Z_{t+1}})\leq\eta\cdot\Delta^{2}$
$\;\;\vdots$
$\text{At step }H:$ $\quad\chi^{2}(P^{(0)}_{Z_{H}}\|P^{(1)}_{Z_{H}})\leq\eta^{H-t}\cdot\Delta^{2}.$
After HtH-t transitions, the divergence shrinks by factor $\eta^{H-t}$. When $\eta<1$, this decay is exponential: the outcome distributions under $H_{0}$ and $H_{1}$ become nearly identical, requiring exponentially many samples to distinguish which hypothesis is true. The next subsection makes this precise.
3.3 The Signal Decay Bound
We now state our main result for this section: a lower bound on the number of i.i.d. outcome samples required to distinguish competing hypotheses about early steps. The bound formalizes the intuition that information decays exponentially as it propagates through the sequential chain. We formalize the hypotheses, prove the bound, derive the critical horizon, and show tightness.
Consider two hypotheses $H_{0}$ and $H_{1}$ about step $t$. These hypotheses induce different distributions $P^{(0)}_{Z_{t}}$ and $P^{(1)}_{Z_{t}}$ over the abstract state $Z_{t}$, and consequently different distributions over the binary outcome $R\in\{0,1\}$. We assume the hypotheses agree on everything else: the transition kernels $K_{u}$ for $u\geq t$ and the terminal map $g$ are identical under both. The question is: how many outcome observations are needed to reliably distinguish $H_{0}$ from $H_{1}$?
Definition 3.7 (Hypothesis Model: Localized Distributional Difference)
Fix a step $t\in\{1,\ldots,H\}$. A pair of hypotheses $(H_{0},H_{1})$ differs only at step $t$ if the following generative model holds:
  1. 1.
    Common initial dynamics: Under both hypotheses, the chain $(Z_{0},\ldots,Z_{t-1})$ follows the same distribution determined by the initial state and transition kernels $K_{0},\ldots,K_{t-2}$.
  2. 2.
    Localized divergence: The conditional distribution of $Z_{t}$ given $Z_{t-1}$ differs:
    $H_{i}:\quad Z_{t}\mid Z_{t-1}\sim K_{t-1}^{(i)}(\cdot\mid Z_{t-1}),\quad i\in\{0,1\}.$
    The induced marginal distributions satisfy $\chi^{2}(P^{(0)}_{Z_{t}}\|P^{(1)}_{Z_{t}})=\Delta^{2}>0$.
  3. 3.
    Shared suffix (downstream dynamics): For $u\geq t$, the transition kernels $K_{u}$ and terminal evaluation $g$ are identical under both hypotheses.
In applications: $H_{0}$ corresponds to “step $t$ is valid” (the transition $K_{t-1}^{(0)}$ produces correct states), while $H_{1}$ corresponds to “step $t$ contains an error” (the transition $K_{t-1}^{(1)}$ may produce erroneous states). In manufacturing: $H_{0}$ means stage $t$ operated correctly; $H_{1}$ means a defect was introduced at stage $t$. The shared-suffix condition isolates step $t$ as the sole source of difference: downstream stages process whatever they receive in the same way under both hypotheses, so any difference in the outcome distribution is attributable entirely to step $t$. Errors introduced at step $t$ may be corrected or may cascade through subsequent stages, but the stages themselves operate identically. We index step $t$ as the abstract state $Z_{t}$ immediately after the $t$-th transition, so HtH-t counts the number of transitions remaining to reach the outcome.
Definition 3.8 (Total Testing Error and Distinguishability)
Let $H_{0}$ and $H_{1}$ induce distributions $P^{(0)}$ and $P^{(1)}$ over outcomes. The total testing error, defined as the sum of type-I and type-II error probabilities minimized over all tests, is
$\mathrm{err}^{*}(P^{(0)},P^{(1)}):=\inf_{\psi}\left(\mathbb{P}_{P^{(0)}}(\psi=1)+\mathbb{P}_{P^{(1)}}(\psi=0)\right)=1-d_{\mathrm{TV}}(P^{(0)},P^{(1)}),$ (11)
where the infimum is over all measurable tests $\psi:\mathcal{Y}\to\{0,1\}$ and $\psi=i$ denotes “accept $H_{i}$.” The hypotheses are $(n,\epsilon)$-distinguishable if the total testing error using $n$ i.i.d. samples is at most $\epsilon$.
Thus achieving total error at most $\epsilon$ requires $d_{\mathrm{TV}}(P^{(0)},P^{(1)})\geq 1-\epsilon$.
Theorem 3.9 (Signal Decay Bound)
Let $H_{0}$ and $H_{1}$ be two hypotheses differing only at step $t$, inducing distributions $P^{(0)}_{Z_{t}}$ and $P^{(1)}_{Z_{t}}$ over the abstract state with $P^{(0)}_{Z_{t}}\ll P^{(1)}_{Z_{t}}$ and $\chi^{2}(P^{(0)}_{Z_{t}}\|P^{(1)}_{Z_{t}})=\Delta^{2}<\infty$. Suppose the Markov state abstraction (Definition 3.1) and Assumption 3.6 hold with contraction coefficient $\eta<1$, and the binary outcome is determined by the terminal abstract state: $R=g(Z_{H})\in\{0,1\}$. Then:
  1. 1.
    Divergence decay: The $\chi^{2}$-divergence at the outcome satisfies
    $\chi^{2}(P^{(0)}_{R}\|P^{(1)}_{R})\leq\eta^{H-t}\cdot\Delta^{2}.$ (12)
  2. 2.
    Sample complexity lower bound: For any target error $\epsilon\in(0,1/2)$, in the regime $\eta^{H-t}\Delta^{2}\leq 1$, any test achieving minimax error at most $\epsilon$ requires
    $n\geq\frac{(1-\epsilon)^{2}}{\eta^{H-t}\cdot\Delta^{2}}$ (13)
    i.i.d. outcome samples. When $\eta^{H-t}\Delta^{2}>1$, this lower bound becomes $O(1)$; in this regime, the outcome distributions are well-separated and $O(1)$ samples suffice to achieve any constant error rate.
  3. 3.
    Exponential growth: For fixed initial divergence $\Delta^{2}=\Theta(1)$, the sample complexity grows exponentially in distance to the outcome:
    $n=\Omega\left(\left(\frac{1}{\eta}\right)^{H-t}\right).$ (14)
Inverting Theorem 3.9 yields a horizon formula: given a fixed sample budget $n$, how far back can one reliably evaluate?
Corollary 3.10 (Critical Horizon)
For initial divergence $\Delta^{2}$, sample budget $n$, and any target error $\epsilon\in(0,1/2)$, a necessary condition for achieving minimax testing error at most $\epsilon$ is that the step lies within the $\epsilon$-critical horizon:
$H-t\leq H_{\mathrm{crit}}(\epsilon):=\frac{\ln\!\bigl(n\Delta^{2}/(1-\epsilon)^{2}\bigr)}{\ln(1/\eta)}=\frac{\ln(n\Delta^{2})-2\ln(1-\epsilon)}{\ln(1/\eta)}.$ (15)
Equivalently, steps more than $H_{\mathrm{crit}}(\epsilon)$ away from the outcome cannot be reliably evaluated from outcome labels alone. For intuition, $H_{\mathrm{crit}}\approx\ln(n\Delta^{2})/\ln(1/\eta)$ when $\epsilon$ is a fixed constant.
Example 3.11 (Critical Horizon in Practice)
Consider contraction coefficient $\eta=0.9$, initial divergence $\Delta^{2}=0.1$, and sample budget $n=10^{6}$. The critical horizon is:
$H_{\mathrm{crit}}=\frac{\ln(10^{6}\times 0.1)}{\ln(1/0.9)}=\frac{\ln(10^{5})}{\ln(1.111)}\approx\frac{11.5}{0.105}\approx 110.$ (16)
With one million outcome samples, we can reliably evaluate steps up to 110 steps before the outcome. Steps further back are beyond the reach of any test.
With stronger contraction $\eta=0.7$ (more diverse sampling):
$H_{\mathrm{crit}}=\frac{11.5}{\ln(1/0.7)}\approx\frac{11.5}{0.357}\approx 32.$ (17)
Stronger mixing dramatically reduces the horizon over which credit assignment is possible. The same sample budget now reaches only 32 steps back. The formula depends logarithmically on $n$ and $\Delta^{2}$ but much more strongly on $\eta$: doubling the sample budget extends the horizon by $\ln 2/\ln(1/\eta)$ steps (about $6.6$ when $\eta=0.9$), while reducing $\eta$ from $0.9$ to $0.7$ cuts it by a factor of three.
Is the exponential rate tight? The following proposition constructs a Bernoulli instance where the likelihood ratio test matches the lower bound up to constants.
Proposition 3.12 (Achievability: Matching Upper Bound)
Consider a stylized two-hypothesis problem where the outcome $R\in\{0,1\}$ is Bernoulli with parameters:
  • Under $H_{0}$: $\mathbb{P}(R=1)=p_{0}$
  • Under $H_{1}$: $\mathbb{P}(R=1)=p_{1}=p_{0}+\delta$ where $\delta=\Theta(\eta^{(H-t)/2}\Delta)$
Then the $\chi^{2}$-divergence between outcome distributions is $\chi^{2}(P^{(0)}_{R}\|P^{(1)}_{R})=\Theta(\eta^{H-t}\Delta^{2})$, and the standard likelihood ratio test achieves minimax error $\epsilon$ with sample complexity $n=O(1/(\eta^{H-t}\Delta^{2}))$.
Together, Theorem 3.9 and Proposition 3.12 show that the exponential rate $(1/\eta)^{H-t}$ is both necessary and sufficient: the Signal Decay Bound is tight up to constants. The exponential barrier is therefore a property of the problem, not of our proof technique.
The preceding results assume exact lumpability (Definition 3.1), but in practice this may hold only approximately. Proposition 10.7 in Appendix 10 shows that if each step introduces at most $\delta$ total-variation error relative to the lumpable idealization, the total-variation distance between outcome distributions acquires an additive penalty of $2(H-t)\delta$ on top of the signal-decay term $\sqrt{\eta^{H-t}\Delta^{2}/2}$. For a 50-stage process with per-step discrepancy $\delta=0.001$, the accumulated penalty is at most $2\times 50\times 0.001=0.1$ in total variation distance. Approximate lumpability becomes a concern only for very deep processes or coarse abstractions, in which case the partition should be refined.
Theorem 3.9 tells us how many samples are needed for reliable testing, and Corollary 3.10 inverts this into a horizon formula. But neither quantifies how bad the error is when the sample budget falls short. The following corollary fills this gap: it gives an explicit lower bound on the minimax testing error, showing that when the available samples $n$ fall below the threshold $1/(\eta^{H-t}\Delta^{2})$ required by Theorem 3.9, the error remains close to $1/2$; no test can do meaningfully better than random guessing.
Corollary 3.13 (Statistical Limits of Value Estimation)
Fix $\epsilon\in(0,1/2)$ and let $R^{1:n}$ denote $n$ i.i.d. outcome samples. Under the Markov state abstraction (Definition 3.1) and Assumption 3.6, the minimax error for testing $H_{0}$ vs. $H_{1}$ based only on $R^{1:n}$ satisfies
$\inf_{\psi}\max_{i\in\{0,1\}}\mathbb{P}_{i}(\psi(R^{1:n})\neq i)\;\geq\;\frac{1}{2}\left(1-\sqrt{\frac{(1+\eta^{H-t}\Delta^{2})^{n}-1}{2}}\right).$
In particular, if
$n\;\leq\;\frac{\ln\!\left(1+2(1-2\epsilon)^{2}\right)}{\eta^{H-t}\Delta^{2}},$
then every test $\psi$ has minimax error at least $\epsilon$.
When $n\eta^{H-t}\Delta^{2}$ is small, $(1+\eta^{H-t}\Delta^{2})^{n}\approx 1+n\eta^{H-t}\Delta^{2}$, so the minimax error is close to $1/2$: the best possible test is barely better than a coin flip. For reinforcement learning, this means that critics trained from outcome labels cannot reliably evaluate steps beyond the critical horizon.
A natural question is whether parallel sampling, generating multiple independent trajectories from the same starting state, can overcome the exponential barrier. The next section shows that it cannot: width provides variance reduction but cannot amplify a signal that has already decayed.
4 Limits of Parallel Sampling
The Signal Decay Bound reveals that outcome signals weaken exponentially with depth. A natural response is to generate more samples: $W$ independent observations from the same starting condition, as in replicated testing or width-based methods like GRPO (DeepSeek-AI 2024). The intuition is that even if individual observations carry weak signal, averaging enough of them should recover the truth. This approach appears to extract more signal from the outcome data that Section 3 proved is fundamentally limited.
But does width truly overcome the exponential barrier, or merely work around it? This section shows that it does not: width reduces noise around an estimate but cannot amplify a signal that has already decayed. We proceed in three steps: derive variance reduction bounds under independence (Section 4.1), show that correlation limits effective width (Section 4.2), and synthesize practical guidance (Section 4.3). Proofs of all results in this section appear in Appendix 11.
4.1 Multi-Rollout Value Estimation
Consider a learner who wishes to estimate the value $V^{\pi}(s):=\mathbb{P}_{\pi}(R=1\mid S_{t}=s)$, the probability of eventual success from state $s$ under the current policy $\pi$. (When conditioning on abstract state, replace $S_{t}$ with $Z_{t}$.) Rather than observing a single trajectory, the learner generates $W$ independent continuations from $s$ and observes their outcomes.
Throughout this section, we analyze the conditional distribution given a realized $s$; unconditional properties follow by averaging over the distribution of $s$.
Definition 4.1 (Multi-Rollout Estimator)
Given a state $s$ and $W$ rollout continuations $\tau_{1},\ldots,\tau_{W}$ from $s$ under policy $\pi$, the multi-rollout value estimator is:
$\hat{V}_{W}(s):=\frac{1}{W}\sum_{j=1}^{W}R_{j},$ (18)
where $R_{j}\in\{0,1\}$ is the binary outcome of rollout $\tau_{j}$.
The estimator is simply the empirical success rate across rollouts. Its statistical properties depend on the independence structure.
{assumption}
[Conditional Independence] Given the conditioning variable $s$ and a fixed policy $\pi$, the rollouts $\tau_{1},\ldots,\tau_{W}$ are conditionally independent:
$P(\tau_{1},\ldots,\tau_{W}\mid s,\pi)=\prod_{j=1}^{W}P(\tau_{j}\mid s,\pi).$ (19)
This assumption holds when each rollout is generated by independent sampling from the policy starting at $s$. It fails when rollouts share randomness, when the policy adapts based on previous rollouts, or when systematic biases cause rollouts to follow similar paths.
Proposition 4.2 (Variance Reduction from Width)
Under Assumption 4.1, the multi-rollout estimator satisfies:
  1. 1.
    Unbiasedness: $\mathbb{E}[\hat{V}_{W}(s)\mid s]=V^{\pi}(s)$.
  2. 2.
    Variance reduction: $\mathrm{Var}(\hat{V}_{W}(s)\mid s)=\dfrac{V^{\pi}(s)(1-V^{\pi}(s))}{W}$.
  3. 3.
    Concentration (Hoeffding): With probability at least $1-\delta$:
    $|\hat{V}_{W}(s)-V^{\pi}(s)|\leq\sqrt{\frac{\ln(2/\delta)}{2W}}.$ (20)
The variance in Part 2 is largest when $V^{\pi}(s)=1/2$ (maximum uncertainty) and zero when outcomes are deterministic. Dividing by $W$ confirms the standard result: to halve the standard deviation, quadruple $W$. Part 3 follows from Hoeffding’s inequality.
Example 4.3 (Width Requirements for Accurate Estimation)
To illustrate the concentration bound, suppose we want to estimate $V^{\pi}(s)$ to within $\pm 0.05$ with 95% confidence. Setting $\delta=0.05$ in Part 3 (so the bound holds with probability $1-\delta=0.95$) gives $\ln(2/\delta)=\ln(40)$. Requiring the error bound to be at most $0.05$:
$\sqrt{\frac{\ln(40)}{2W}}\leq 0.05\quad\Longrightarrow\quad W\geq\frac{\ln(40)}{2\times 0.0025}\approx 738.$ (21)
Approximately 740 rollouts per state suffice. This is a worst-case bound; when $V^{\pi}(s)$ is near 0 or 1, variance is smaller and fewer rollouts suffice. Relaxing to $\pm 0.1$ accuracy reduces the requirement to about 185 rollouts.
The preceding analysis assumes independent rollouts. In practice, rollouts from the same starting point are correlated, which further limits the benefit of width.
4.2 Correlation Limits Effective Width
To quantify this limitation, we must account for correlation among rollouts from the same prefix. Consider collecting data from $n$ independent starting points, generating $W$ rollouts from each. Rollouts from different starting points are independent, and their outcomes are uncorrelated. However, rollouts from the same starting point share the same prefix and tend to make similar errors, inducing positive correlation among their outcomes. Correlation arises from three principal mechanisms:
  1. 1.
    Shared prefix errors: When a prefix contains a latent mistake, all continuations inherit the defect. If an early assembly stage installs a component off-center, every subsequent inspection reveals the same misalignment; if a language model makes an algebraic error in step 3, all continuations from that point inherit it.
  2. 2.
    Mode collapse: Low-entropy policies generate nearly identical continuations from the same prefix ($\rho\to 1$). Low-temperature sampling may produce only 2–3 distinct reasoning paths from $W=50$ rollouts; a deterministic assembly line yields essentially identical outputs from the same inputs.
  3. 3.
    Systematic biases: Learned patterns produce consistent error modes activated by specific inputs. A model that consistently confuses correlation with causation will make the same inferential error across all rollouts from a given prefix, regardless of token-level variation.
When $W=1$ (one rollout per starting point), we obtain $n$ i.i.d. outcome observations. When $W>1$, this within-starting-point correlation reduces the effective sample size, as the following theorem quantifies.
Theorem 4.4 (Effective Width Under Correlation)
Consider $W$ rollouts from a single starting point $s$. Assume the outcomes satisfy equicorrelation: $\mathrm{Corr}(R_{i},R_{j}\mid s)=\rho$ for all $i\neq j$, where $\rho\in[0,1)$ is a constant. (More generally, let $\rho$ be an upper bound on pairwise correlations.) The variance of the multi-rollout estimator $\hat{V}_{W}(s)=\frac{1}{W}\sum_{j=1}^{W}R_{j}$ satisfies:
$\mathrm{Var}(\hat{V}_{W}(s)\mid s)=\frac{\sigma^{2}(s)}{W}\cdot(1+(W-1)\rho),$ (22)
where $\sigma^{2}(s)=V^{\pi}(s)(1-V^{\pi}(s))$. Equivalently, the correlated estimator has the same variance as an estimator using $W_{\mathrm{eff}}$ independent rollouts, where the effective width is:
$W_{\mathrm{eff}}:=\frac{W}{1+(W-1)\rho}.$ (23)
When $\rho=0$ (perfect independence), $W_{\mathrm{eff}}=W$ and no degradation occurs. When $\rho>0$, the effective width saturates as $W\to\infty$: $W_{\mathrm{eff}}\to 1/\rho$.
Example 4.5 (Impact of Correlation)
Suppose rollouts have correlation $\rho=0.2$ (modest correlation from shared reasoning patterns). The effective width for various nominal widths:
$W$ 10 50 100 500 $\infty$
$W_{\mathrm{eff}}$ 3.6 4.6 4.8 5.0 5.0
Even with 500 rollouts, the effective width is only 5. Beyond roughly $1/\rho=5$ rollouts, additional sampling provides negligible benefit.
Theorem 4.4 shows that correlation reduces effective width from $W$ to $W_{\mathrm{eff}}$. How does this affect the critical horizon? While the $W$ rollouts within each starting point are correlated, the $n$ starting points are independent. Define the effective sample size as $N_{\mathrm{eff}}:=n\cdot W_{\mathrm{eff}}$. Combined with the Signal Decay Bound (Theorem 3.9), distinguishing hypotheses about step $t$ requires
$N_{\mathrm{eff}}\;\gtrsim\;\frac{1}{\eta^{H-t}\Delta^{2}}.$ (24)
The following corollary formalizes the resulting critical horizon.
Corollary 4.6 (Critical Horizon with Width)
Let $n$ be the number of independent starting points and $W_{\mathrm{eff}}$ the effective width per starting point. The critical horizon extends to
$H_{\mathrm{crit}}=\frac{\ln(n\cdot W_{\mathrm{eff}}\cdot\Delta^{2})}{\ln(1/\eta)}.$
Width enters the critical horizon only logarithmically: doubling $W_{\mathrm{eff}}$ adds merely $\ln(2)/\ln(1/\eta)\approx 6.6$ steps when $\eta=0.9$. And correlation caps the benefit: since $W_{\mathrm{eff}}\leq 1/\rho$, the effective sample size cannot exceed $n/\rho$ regardless of nominal width. The exponential dependence on depth cannot be escaped by increasing $W$.
4.3 Summary: When Width Works
The preceding analysis yields two necessary conditions for width to be beneficial:
  1. 1.
    Signal must exist: The step must lie within the critical horizon ($H-t\leq H_{\mathrm{crit}}$). Beyond this, outcome data is uninformative regardless of sample size.
  2. 2.
    Rollouts must be approximately independent: Effective width saturates at $1/\rho$, so high correlation eliminates the benefit of additional rollouts.
When either condition fails, width provides diminishing returns. The following decision rule provides a practical roadmap for determining when width-based methods can succeed:
Practical Decision Rule for Width-Based Methods 1. Estimate correlation $\rho$ from pilot data or domain knowledge (e.g., deterministic policies have $\rho\approx 1$; high-temperature sampling has lower $\rho$). Compute $W_{\mathrm{eff}}=W/(1+(W-1)\rho)$. 2. Compute effective horizon: $H_{\mathrm{crit}}=\ln(n\cdot W_{\mathrm{eff}}\cdot\Delta^{2})/\ln(1/\eta)$, where $\Delta^{2}$ is the initial $\chi^{2}$-divergence between hypotheses and $\eta$ is the contraction coefficient. 3. Decision: If $H\leq H_{\mathrm{crit}}$, width-based methods can succeed. Otherwise, process supervision is necessary.
The next section addresses how step-level supervision, once obtained, should be aggregated into training objectives.
5 From Outcome Limits to Process Supervision
The preceding sections established that outcome-based credit assignment faces an exponential barrier (Theorem 3.9) that width can only logarithmically mitigate (Section 4). For deep processes where $H>H_{\mathrm{crit}}$, this leaves only one path forward: step-level supervision. Process reward models (Lightman et al. 2024), intermediate checkpoints in manufacturing, and stage-gate reviews in project management all implement this principle. Step-level labels are not merely convenient; they are necessary.
But step-level supervision introduces its own challenge: how should step-level feedback be aggregated into a training or evaluation objective? Standard reinforcement learning sums rewards across steps (Sutton et al. 2000), but for sequential validity, where a single error invalidates the entire chain, this additive aggregation optimizes the wrong quantity. Proofs of all results in this section appear in Appendix 12.
5.1 The Objective Mismatch
Given that step-level supervision is necessary for deep processes, how should it be used? When training a system with step-level labels $r_{t}\in\{0,1\}$ indicating whether each step is valid, one can maximize either the expected number of correct steps or the probability that all steps are correct. These correspond, respectively, to additive and multiplicative aggregation of step-level feedback.
Definition 5.1 (Step-Quality and Validity Objectives)
For a policy $\pi$ generating trajectories of length $H$ from initial state $S_{0}$, with step validity indicators $r_{0},\ldots,r_{H-1}$:
  1. 1.
    The step-quality objective is $J_{\mathrm{add}}:=\mathbb{E}_{\pi}\left[\sum_{t=0}^{H-1}r_{t}\,\big|\,S_{0}\right]$, the expected number of correct steps.
  2. 2.
    The validity objective is $J_{\mathrm{mult}}:=\mathbb{E}_{\pi}\left[\prod_{t=0}^{H-1}r_{t}\,\big|\,S_{0}\right]$, the probability that all steps are correct.
Multiplicative objectives appear in risk-sensitive control (Howard and Matheson 1972) and stochastic reachability (Abate et al. 2008). Our contribution is identifying that sequential validity, where a single failure invalidates the entire chain, naturally requires this structure, and analyzing the optimization challenges it creates. This applies to logical reasoning, manufacturing processes where any defective stage produces a defective product, and service operations where any failed touchpoint loses the customer.
Under independent steps with common success probability $p$, we have $J_{\mathrm{add}}=Hp$ while $J_{\mathrm{mult}}=p^{H}$. Even at 99% per-step accuracy over 100 steps, $J_{\mathrm{add}}=99$ while $J_{\mathrm{mult}}=0.37$: systems optimizing additive objectives learn to produce sequences that are mostly correct but nonetheless invalid.
5.2 Gradient Decay in the Multiplicative Objective
The preceding analysis suggests using the multiplicative objective directly. However, this approach encounters a fundamental optimization barrier: the multiplicative objective exhibits vanishing gradients for long chains with imperfect per-step accuracy.
Under the independence model, differentiating $J_{\mathrm{add}}=Hp$ and $J_{\mathrm{mult}}=p^{H}$ with respect to policy parameters $\theta$ gives $\nabla_{\theta}J_{\mathrm{add}}=H\nabla_{\theta}p$ and $\nabla_{\theta}J_{\mathrm{mult}}=Hp^{H-1}\nabla_{\theta}p$. The multiplicative gradient contains an additional factor $p^{H-1}$ that attenuates the learning signal. For $p=0.95$ over a 100-step chain, this factor equals $0.95^{99}\approx 0.006$, rendering the validity objective nearly unlearnable. The barrier becomes severe precisely when it matters most: for longer chains and policies that have not yet achieved high per-step accuracy.
This gradient decay is conceptually distinct from the signal decay of Section 3. Signal decay concerns information flow: outcome data becomes uninformative about early steps. Gradient decay concerns objective structure: the multiplicative objective attenuates learning signal regardless of information availability. A complete training framework must address both barriers. We propose a curriculum approach: begin with the learnable additive objective, then transition to the correct multiplicative objective as policy quality improves (Bengio et al. 2009).
Definition 5.2 (Interpolated Objective)
For $\alpha\in[0,1]$, the interpolated objective is:
$J_{\alpha}(\theta):=(1-\alpha)\cdot J_{\mathrm{add}}(\theta)+\alpha\cdot J_{\mathrm{mult}}(\theta).$ (25)
At $\alpha=0$, the objective is learnable but incorrect; at $\alpha=1$, correct but potentially unlearnable. The following result characterizes how $\alpha$ should evolve with policy quality.
Proposition 5.3 (Gradient-Preserving Curriculum)
Assume $\nabla_{\theta}J_{\mathrm{add}}$ and $\nabla_{\theta}J_{\mathrm{mult}}$ have non-negative inner product. To maintain $\|\nabla_{\theta}J_{\alpha}\|\geq c\cdot\|\nabla_{\theta}J_{\mathrm{add}}\|$ for some $c\in(0,1)$, it suffices to set $\alpha\leq 1-c$.
The proposition establishes that the interpolated objective preserves gradient signal when $\alpha$ is sufficiently small. In the independence model where $J_{\mathrm{add}}=Hp$ and $J_{\mathrm{mult}}=p^{H}$, the gradients satisfy $\|\nabla_{\theta}J_{\mathrm{mult}}\|=p^{H-1}\|\nabla_{\theta}J_{\mathrm{add}}\|$. As the policy improves and per-step success probability $p$ approaches $1$, this attenuation factor approaches $1$, allowing $\alpha$ to increase toward $1$ without sacrificing gradient magnitude. This yields a curriculum: begin training with $\alpha$ near 0 to ensure learnability, then gradually increase $\alpha$ toward $1$ as the policy improves, transitioning from a learnable but misspecified objective to one that is correctly specified.
Having established why step-level supervision is necessary and how to aggregate it, the next section addresses where: given an inspection budget, which steps should be observed?
6 Optimal Intermediate Inspection Design
We now convert the information-theoretic limits into optimal intermediate inspection schedules. The organizing principle is that inspections shorten the information path that credit assignment must traverse: instead of propagating from step $t$ to the terminal outcome at time $H$, it suffices to propagate from $t$ to the next checkpoint.
The inspection design problem has a long history in OR under different guises. In manufacturing, it is the placement of quality checkpoints along a production line (Lindsay and Bishop 1964, Raz 1986). In sequential testing, it is related to adaptive group testing (Du and Hwang 2000). Our contribution is to solve inspection placement under the specific signal-decay structure implied by the Strong Data Processing Inequality, yielding closed-form schedules for both homogeneous and heterogeneous contraction.
We proceed in three steps. First, we formalize the inspection feedback model and extend the signal-decay bound to an arbitrary inspection time $u$. Second, under homogeneous contraction we show that worst-case sample complexity is governed by a single scalar, the largest distance from any step to its next inspection, and we solve the resulting placement problem (uniform spacing). Third, under heterogeneous contraction we show that feasibility depends on cumulative information loss along each segment, and we solve the minimum-inspection design problem via a greedy information-distance schedule. Cost and sample tradeoffs are then handled in Section 6.5. Proofs of all results in this section appear in Appendix 13.
6.1 Inspection Schedules and Feedback Model
Recall the abstract Markov chain $(Z_{0},\dots,Z_{H})$ induced by the abstraction $\phi$ and policy $\pi$. We treat the terminal outcome $R=g(Z_{H})\in\{0,1\}$ as a terminal inspection at time $H$.
Definition 6.1 (Inspection Schedule)
An inspection schedule is a strictly increasing set of inspection times
$\mathcal{T}=\{t_{1}<t_{2}<\cdots<t_{m}\}\subseteq\{1,2,\dots,H-1\}.$
We use the augmented sequence $0=:t_{0}<t_{1}<\cdots<t_{m}<t_{m+1}:=H$. The number of paid intermediate inspections is $m$.
Definition 6.2 (Inspection Observations)
At each inspection time $u\in\mathcal{T}\cup\{H\}$, the learner observes an inspection output $Y_{u}:=g_{u}(Z_{u})\in\mathcal{Y}_{u}$, where $g_{u}$ is an external evaluator (human, verifier, metrology device) applied to the abstract state. The terminal outcome is the special case $Y_{H}\equiv R=g(Z_{H})$.
The model allows many inspection types: prefix-validity checks, milestone verification, partial correctness scores, or any measurement that is a function of the abstract state. The results below are information-theoretic and do not assume any particular learning algorithm.
6.2 Signal Decay to Intermediate Inspections
The design lever is that intermediate checkpoints reduce the effective distance over which information must propagate. If step $t$ is followed by an inspection at time $u<H$, then credit assignment for step $t$ need only overcome utu-t steps of attenuation rather than HtH-t. We formalize this for inhomogeneous contraction (step-dependent mixing rates); the homogeneous case follows in Section 6.3.
{assumption}
[Inhomogeneous Contractive Dynamics] Let $K_{t}$ denote the abstract transition kernel from $Z_{t}$ to $Z_{t+1}$, with $\chi^{2}$-contraction coefficient $\eta_{\chi^{2}}(K_{t})\leq\eta_{t}\in(0,1]$ for $t=0,1,\ldots,H-1$. Define the cumulative attenuation from step $t$ to inspection time $u>t$ as
$\mathrm{Attn}(t\to u):=\prod_{j=t}^{u-1}\eta_{j}.$
In the homogeneous case $\eta_{t}\equiv\eta$, this simplifies to $\mathrm{Attn}(t\to u)=\eta^{u-t}$.
Proposition 6.3 (Signal Decay to an Intermediate Inspection)
Consider two hypotheses $H_{0}$ and $H_{1}$ differing only at step $t$, inducing abstract state distributions with $\chi^{2}(P^{(0)}_{Z_{t}}\|P^{(1)}_{Z_{t}})=\Delta^{2}$. Under Assumption 6.2, the $\chi^{2}$-divergence at any downstream inspection $Y_{u}=g_{u}(Z_{u})$ satisfies
$\chi^{2}(P^{(0)}_{Y_{u}}\|P^{(1)}_{Y_{u}})\leq\mathrm{Attn}(t\to u)\cdot\Delta^{2}.$
Proposition 6.3 implies that an inspection at time $u$ replaces the terminal attenuation factor $\mathrm{Attn}(t\to H)$ by the shorter-horizon factor $\mathrm{Attn}(t\to u)$.
Corollary 6.4 (Sample Complexity with Intermediate Inspection)
Under the conditions of Proposition 6.3, any test distinguishing $H_{0}$ from $H_{1}$ at error rate $\epsilon$ requires
$n\geq\frac{(1-\epsilon)^{2}}{\mathrm{Attn}(t\to u)\cdot\Delta^{2}}.$
In the homogeneous case, $n\geq(1-\epsilon)^{2}/(\eta^{u-t}\Delta^{2})$.
Corollary 6.4 makes the value of intermediate inspection precise: placing a checkpoint at time $u$ reduces the attenuation exponent from HtH-t to utu-t. The remaining question is how to place inspection times so that no step is too far from its next checkpoint.
6.3 Homogeneous Contraction: Minimax-Optimal Placement
Under homogeneous contraction, the sample complexity penalty depends only on the distance from a step to its next inspection (Corollary 6.4). This turns inspection design into a one-dimensional placement problem on $\{0,1,\ldots,H\}$ with a single governing design statistic.
The development in this subsection follows a short chain of implications. We first define the distance of each step to its next checkpoint and the induced maximal gap $L(\mathcal{T})$. We then show that any schedule must incur a worst-case sample complexity that depends exponentially on $L(\mathcal{T})$. Finally, we solve the placement problem by minimizing $L(\mathcal{T})$ over schedules with a fixed number of inspections $m$ and translate the result into a minimum number of inspections required for feasibility.
Definition 6.5 (Downstream Distance and Maximal Gap)
For a schedule $\mathcal{T}=\{t_{1}<\cdots<t_{m}\}$ with augmented times $t_{0}=0$ and $t_{m+1}=H$, the downstream distance of step $t$ is the distance to its next inspection:
$d_{\mathcal{T}}(t):=\min\{u-t:u\in\mathcal{T}\cup\{H\},\ u>t\}.$
The maximal gap is $L(\mathcal{T}):=\max_{t\in\{0,\ldots,H-1\}}d_{\mathcal{T}}(t)$, equivalently $L(\mathcal{T})=\max_{i=0,\ldots,m}(t_{i+1}-t_{i})$.
The maximal gap determines worst-case performance: the step at the beginning of the longest segment between inspections faces the most attenuation before any observation. For a fixed number of intermediate inspections $m$, the minimax placement problem is therefore
$\min_{|\mathcal{T}|=m}L(\mathcal{T}),$
because $L(\mathcal{T})$ controls the worst-case attenuation exponent (and hence the worst-case sample complexity) across all steps. Operationally, $L(\mathcal{T})$ is the length of the longest uninspected “run” of steps; the minimax goal is to prevent any stage from being too far from a checkpoint.
Proposition 6.6 (Worst-Case Sample Complexity)
Under homogeneous contraction $\eta_{t}\equiv\eta<1$, any inspection schedule $\mathcal{T}$ admits a step $t^{\star}$ such that distinguishing hypotheses about $t^{\star}$ requires
$n\geq\frac{(1-\epsilon)^{2}}{\eta^{L(\mathcal{T})}\Delta^{2}}$
samples.
Proposition 6.6 identifies the correct minimax objective: reduce the largest gap to reduce worst-case sample complexity. The remaining question is purely combinatorial: how should $m$ inspections partition a horizon of length $H$ to minimize the length of the longest segment?
Theorem 6.7 (Minimax-Optimal Inspection Placement)
Under homogeneous contraction $\eta_{t}\equiv\eta<1$, the minimal achievable maximal gap with $m$ intermediate inspections is
$\min_{|\mathcal{T}|=m}L(\mathcal{T})=\Big\lceil\frac{H}{m+1}\Big\rceil.$
This minimum is achieved by near-uniform spacing: placing inspections at $t_{i}=\lfloor iH/(m+1)\rfloor$ for $i=1,\ldots,m$.
The implication is that the detailed location of checkpoints is secondary; what matters is controlling the longest uninspected run. Combining Proposition 6.6 and Theorem 6.7 yields an explicit feasibility criterion.
Corollary 6.8 (Sample Complexity Under $m$ Inspections)
With $m$ minimax-optimal inspections, the worst-case sample complexity for credit assignment is
$n\geq\frac{(1-\epsilon)^{2}}{\eta^{\lceil H/(m+1)\rceil}\Delta^{2}}.$
Corollary 6.9 (Minimum Inspections for Feasibility)
For all steps to be $(n,\epsilon)$-distinguishable, the number of inspections must satisfy
$m\geq\Big\lceil\frac{H}{H_{\mathrm{crit}}(\epsilon)}\Big\rceil-1,$
where $H_{\mathrm{crit}}(\epsilon)=\bigl(\ln(n\Delta^{2})-2\ln(1-\epsilon)\bigr)/\ln(1/\eta)$ is the $\epsilon$-critical horizon. This is a necessary condition; the sufficient number may be one higher due to integer rounding when $H/(m+1)$ is not an integer.
Example 6.10 (Semiconductor Fabrication)
A semiconductor process has $H=50$ stages with $\eta=0.85$. Terminal testing costs \$10/wafer; intermediate metrology costs \$50/inspection. With n=10,000n=10{,}000 wafers, $\Delta^{2}=0.2$, and target error $\epsilon=0.1$:
$H_{\mathrm{crit}}(\epsilon)=\frac{\ln(2000)-2\ln(0.9)}{\ln(1/0.85)}\approx 48.1<50.$
Since $H>H_{\mathrm{crit}}(\epsilon)$, outcome-only evaluation is insufficient. Corollary 6.9 gives $m\geq\lceil 50/48.1\rceil-1=1$. One inspection at the midpoint ($t_{1}=25$) reduces the maximal gap to 25, bringing all stages within the critical horizon. The inspection adds \$50/wafer but makes the study feasible; without it, the required sample size would exceed the budget.
6.4 Heterogeneous Contraction: Minimum Inspections via Information Distance
Uniform spacing is optimal when all steps contract equally, but many processes exhibit depth-dependent contraction. In such settings, the design objective shifts from “minimize the largest segment length” to a feasibility-driven problem: choose as few inspections as possible so that every segment has cumulative information loss at most the threshold implied by the sample budget.
Definition 6.11 (Information Distance)
Under Assumption 6.2, the per-step information distance is $w_{t}:=\ln(1/\eta_{t})\geq 0$. For an interval $[a,b)$, the cumulative information distance is
$W([a,b)):=\sum_{t=a}^{b-1}w_{t}=\ln(1/\mathrm{Attn}(a\to b)).$
Steps with strong contraction (small $\eta_{t}$) contribute large $w_{t}$; steps with weak contraction contribute little.
Because $\mathrm{Attn}(t\to u)=\exp(-W([t,u)))$, Corollary 6.4 yields a segment-level feasibility threshold.
Lemma 6.12 (Per-Segment Feasibility)
Fix consecutive inspection times $t_{i}<t_{i+1}$. If
$W([t_{i},t_{i+1}))\leq\Gamma(n,\Delta^{2},\epsilon):=\ln(n\Delta^{2})-2\ln(1-\epsilon),$
then every step $t\in\{t_{i},\ldots,t_{i+1}-1\}$ is $(n,\epsilon)$-distinguishable using the inspection at $t_{i+1}$. Conversely, if step $t_{i}$ is $(n,\epsilon)$-distinguishable using the inspection at $t_{i+1}$, then the displayed inequality must hold.
The heterogeneous inspection design problem is to partition the horizon into segments satisfying the Lemma 6.12 constraint while using as few paid inspections as possible. A greedy construction is optimal.
Theorem 6.13 (Greedy Inspection is Optimal)
Among all schedules satisfying $W([t_{i},t_{i+1}))\leq\Gamma$ for each segment, the following greedy construction minimizes the number of inspections: starting at $t_{0}=0$, place each inspection $t_{i+1}$ at the largest index such that $W([t_{i},t_{i+1}))\leq\Gamma$.
Because greedy stretches each segment to its feasibility limit, inspections concentrate automatically in regions where contraction is strongest.
When $\eta_{t}\equiv\eta$, we have $w_{t}\equiv\ln(1/\eta)$, so greedy produces segments of length $\lfloor\Gamma/\ln(1/\eta)\rfloor$, recovering uniform spacing (Theorem 6.7). When contraction varies, greedy produces non-uniform placement: segments spanning high-contraction steps are short; segments spanning low-contraction steps are long.
Example 6.14 (Service Operations)
A customer journey has $H=50$ touchpoints. Early touchpoints (awareness, consideration) have high variability ($\eta_{t}=0.6$ for $t\leq 10$), while later touchpoints are more deterministic ($\eta_{t}=0.95$ for $t>10$). With n=1,000n=1{,}000 customers, $\Delta^{2}=0.3$, and $\epsilon=0.1$, the feasibility threshold is $\Gamma=\ln(300)-2\ln(0.9)\approx 5.9$.
The 11 high-contraction steps contribute $W([0,11))=11\times\ln(1/0.6)\approx 5.6$, nearly exhausting the budget. The greedy algorithm extends through 5 additional low-contraction steps before exceeding $\Gamma$, placing an inspection at $t_{1}=16$ (since $W([0,16))\approx 5.9\leq\Gamma$ but $W([0,17))\approx 5.9>\Gamma$). The remaining 34 low-contraction steps contribute only $W([16,50))=34\times\ln(1/0.95)\approx 1.7<\Gamma$, fitting in one segment. The optimal schedule has one inspection, placed where contraction is strongest, demonstrating that non-uniform placement arises naturally when contraction varies.
6.5 Joint Sample-Inspection Budgeting
The preceding results characterize inspection placement in information units (maximal gap under homogeneous contraction, and information-distance segments under heterogeneous contraction). We now incorporate operational costs: in practice, the number of inspections is itself a design choice subject to a budget constraint. Let $c_{\mathrm{out}}$ denote the cost per trajectory for terminal evaluation and $c_{\mathrm{insp}}$ the cost per intermediate inspection per trajectory. Collecting $n$ trajectories with $m$ inspections each costs $B=n(c_{\mathrm{out}}+m\cdot c_{\mathrm{insp}})$.
Proposition 6.15 (Budget Lower Bound)
Under homogeneous contraction, the minimum budget required for worst-case $(n,\epsilon)$-distinguishability with $m$ inspections is
$B\geq(c_{\mathrm{out}}+m\cdot c_{\mathrm{insp}})\cdot\frac{(1-\epsilon)^{2}}{\eta^{\lceil H/(m+1)\rceil}\Delta^{2}}.$
Proposition 6.15 reveals a tradeoff: increasing $m$ reduces the sample requirement exponentially, but raises per-trajectory cost linearly. A simple rule is to increase $m$ until the maximal gap $\lceil H/(m+1)\rceil$ falls below $H_{\mathrm{crit}}(\epsilon)$, at which point further inspections yield negligible sample savings at the target error level.
Corollary 6.16 (Inspection Density for Polynomial Complexity)
To achieve polynomial sample complexity $n=O(H^{p})$ under homogeneous contraction, the number of inspections must scale as $m=\Omega(H/\log H)$.
Corollary 6.16 implies that inspection effort must grow nearly linearly with process depth to avoid exponential sample requirements. We summarize the preceding results in the following inspection design procedure.
Inspection Design Procedure 1. Estimate contraction: Bound the contraction coefficient $\eta$ (or $\{\eta_{t}\}$) using A/B comparisons, cohort tracking, or variance decay measurements. 2. Compute feasibility threshold: $\Gamma(n,\Delta^{2},\epsilon)=\ln(n\Delta^{2})-2\ln(1-\epsilon)$. Equivalently, under homogeneous contraction compute $H_{\mathrm{crit}}(\epsilon)=\Gamma/\ln(1/\eta)$. 3. Choose inspections for feasibility: Under homogeneous contraction, choose the smallest $m$ satisfying $m\geq\lceil H/H_{\mathrm{crit}}(\epsilon)\rceil-1$. Under heterogeneous contraction, choose the smallest $m$ produced by the greedy rule in Theorem 6.13. 4. Place inspections: Uniformly at $t_{i}=\lfloor iH/(m+1)\rfloor$ under homogeneous contraction (Theorem 6.7), or via greedy information-distance placement under heterogeneous contraction (Theorem 6.13). 5. Budget the study: Set $n$ and $m$ to satisfy $B=n(c_{\mathrm{out}}+mc_{\mathrm{insp}})$ and compare against the lower bound in Proposition 6.15.
The theoretical results yield three practical insights. First, intermediate inspection is often an investment in feasibility, not merely precision: a single well-placed checkpoint can make an otherwise impossible attribution problem tractable (Example 6.10). Second, uniform spacing provides a robust default when contraction rates are uncertain (Theorem 6.7); when stage-specific estimates are available, the greedy algorithm (Theorem 6.13) concentrates inspections where contraction is strongest. Third, inspection effort must scale with process depth: organizations extending their value chains cannot hold inspection budgets constant (Corollary 6.16).
7 Numerical Validation
We validate five core theoretical predictions using synthetic Markov chains with controlled parameters. For effective width and objective mismatch, we additionally demonstrate the predicted phenomena on LLM reasoning chains from GSM8K. Synthetic experiments allow exact control of the contraction coefficient $\eta$, providing clean verification of the bounds; LLM experiments confirm that the phenomena occur in practice. The experiments cover signal decay (Section 7.1), effective width under correlation (Section 7.2), optimal inspection design (Section 7.3), the critical horizon phase transition (Section 7.4), and the objective mismatch between additive and multiplicative rewards (Section 7.5). All code will be made available upon acceptance.
7.1 Signal Decay Bound
Theorem 3.9 establishes that $\chi^{2}$-divergence between outcome distributions decays exponentially with distance from outcome:
$\chi^{2}(P_{S_{H}}\|Q_{S_{H}})\;\leq\;\eta^{H-t}\cdot\chi^{2}(P_{S_{t}}\|Q_{S_{t}}).$
To test this bound, we construct synthetic Markov chains with 10 abstract states and exact contraction coefficient $\eta\in\{0.7,0.8,0.9,0.95\}$, using the kernel $K=\sqrt{\eta}\,I+(1-\sqrt{\eta})\,(1/10)\,\mathbf{1}\mathbf{1}^{\top}$, a convex combination of the identity and the uniform-mixing matrix whose $\chi^{2}$-contraction coefficient is exactly $\eta$. For each step $t$, we set $P_{t}=\delta_{0}$ (point mass on state 0) and $Q_{t}=\pi$ (the uniform stationary distribution), giving $\chi^{2}(P_{t}\|Q_{t})=9$. We propagate both distributions to the terminal state through the contractive kernel and measure the resulting $\chi^{2}$-divergence at the state distribution level.
Figure 7.1 plots the measured divergence against distance to outcome (Ht)(H-t) on a logarithmic scale for $H=40$, showing all four values of $\eta$. The dashed lines show the theoretical prediction $\eta^{H-t}\cdot\Delta^{2}$ with $\Delta^{2}=9$. Because the reference distribution $Q_{t}$ equals the stationary distribution $\pi$, the bound is achieved with equality: measured divergences lie exactly on the theoretical lines. The measurements follow perfect linear trends on the log scale with distinct slopes for each $\eta$, confirming exponential decay at the predicted rates. Results for other horizons ($H\in\{10,20\}$) are identical in pattern.
[…]
Validation of the Signal Decay Bound (Theorem 3.9). Measured $\chi^{2}$-divergences (dots) match the theoretical prediction $\eta^{H-t}\cdot\Delta^{2}$ (dashed lines) exactly for all four contraction rates, confirming exponential decay with rate $\eta$. Setup: 10-state Markov chains with $H=40$ and $\Delta^{2}=9$.
7.2 Effective Width Under Correlation
Theorem 4.4 shows that positively correlated rollouts yield diminished variance reduction:
$W_{\mathrm{eff}}\;=\;\frac{W}{1+(W-1)\rho},$
where $\rho$ is the pairwise correlation between rollout outcomes. As $W\to\infty$, effective width saturates at $1/\rho$, meaning that beyond a certain point, additional rollouts provide negligible benefit.
We test this prediction in two settings. In synthetic experiments, we induce correlation $\rho=0.15$ by adding a shared noise component across rollouts, then measure variance reduction at $W\in\{1,4,16,64,256\}$. In LLM experiments, we generate $W\in\{1,4,16\}$ rollouts per GSM8K problem and measure the empirical pairwise correlation ($\bar{\rho}\approx 0.49$, averaged across $W\in\{4,16\}$) and variance reduction.
Figure 7.2 shows the results. In both panels, measured effective width (red dots) closely tracks the theoretical prediction (blue curve). The synthetic experiment confirms saturation: at $W=256$ rollouts, effective width reaches only $6.6$, compared to the theoretical maximum of $1/\rho=6.7$. The LLM experiment shows even stronger saturation due to higher correlation: $W=16$ nominal rollouts yield effective width of only $2.2$.
[…]
Validation of the Effective Width Formula (Theorem 4.4). Measured effective width (dots) matches the theoretical prediction $W/(1+(W-1)\rho)$ (curve) and saturates at $1/\rho$. (a) Synthetic with $\rho=0.15$. (b) LLM with $\rho\approx 0.49$.
7.3 Optimal Inspection Design
Theorem 6.7 establishes that under homogeneous contraction, uniformly spaced checkpoints minimize the worst-case attribution error across all steps. The intuition is that any gap between checkpoints allows signal to decay; uniform spacing ensures no gap is larger than necessary.
We test this with $H=20$ steps, contraction coefficient $\eta=0.9$, and $m=3$ intermediate checkpoints, comparing four schedules: uniform placement at $\{5,10,15\}$, front-loaded at $\{2,4,6\}$, back-loaded at $\{14,16,18\}$, and random at $\{2,13,14\}$. For each schedule, we compute the worst-case attribution error across all 20 steps.
Figure 7.3 shows the results. Uniform spacing achieves worst-case error of 0.41, while the alternatives achieve 0.69 (random) to 0.77 (front- and back-loaded). This represents a 40–47% reduction in worst-case error, confirming that uniform spacing is not merely a reasonable heuristic but the optimal choice under the conditions of the theorem.
[…]
Validation of Optimal Inspection Design (Theorem 6.7). Uniform checkpoint spacing (green) achieves 40–47% lower worst-case error than front-loaded, back-loaded, or random schedules. Checkpoint positions shown above each bar.
7.4 Critical Horizon Phase Transition
Corollary 3.10 predicts a sharp phase transition: steps within $H_{\mathrm{crit}}\approx\ln(n\Delta^{2})/\ln(1/\eta)$ of the outcome can be reliably attributed, while steps beyond this horizon cannot. We test this prediction by measuring attribution accuracy as a function of distance from outcome.
We use $H=40$ steps, 10 abstract states, n=1,000n=1{,}000 samples per hypothesis, and $\Delta^{2}=9$. We test two contraction rates: $\eta=0.7$ (predicted $H_{\mathrm{crit}}\approx 26$) and $\eta=0.8$ (predicted $H_{\mathrm{crit}}\approx 41$). For each step position, we generate outcome samples under two initial distributions and test whether the step’s identity can be recovered.
Figure 7.4 shows the results. Attribution accuracy for $\eta=0.7$ drops from 89% at distance 1 to chance level (50%) around distance 23, slightly before the theoretical $H_{\mathrm{crit}}\approx 26$. For $\eta=0.8$, the slower decay rate places $H_{\mathrm{crit}}\approx 41$ beyond the maximum horizon $H=40$, so accuracy remains above chance throughout but declines steadily. The critical horizon formula provides a tight upper bound on the reach of outcome-based credit assignment: accuracy degrades continuously with distance and reaches chance level at or before $H_{\mathrm{crit}}$, confirming that steps beyond the critical horizon cannot be reliably evaluated. The gap between the empirical transition and the theoretical bound reflects the difference between the simple classifier used here and the information-theoretically optimal test.
[…]
Validation of the Critical Horizon (Corollary 3.10). Attribution accuracy drops to chance (50%) at or before the predicted $H_{\mathrm{crit}}$ (dashed lines) for both contraction rates. For $\eta=0.7$, $H_{\mathrm{crit}}\approx 26$; for $\eta=0.8$, $H_{\mathrm{crit}}\approx 41$, which exceeds the horizon $H=40$, so accuracy remains above chance throughout.
7.5 Objective Mismatch
Definition 5.1 identifies a fundamental gap between step-quality ($J_{\mathrm{add}}$) and validity ($J_{\mathrm{mult}}$) objectives: a chain can have most steps correct yet still produce a wrong answer. We test whether this “mostly correct fallacy” occurs in practice by evaluating 50 GSM8K reasoning chains, scoring each step for correctness and comparing additive rewards (fraction of steps rated correct) against a binary validity indicator (1 if the final answer is correct, 0 otherwise).
Figure 7.5(a) plots each chain’s additive reward (x-axis) against its validity indicator (y-axis: 1 for correct, 0 for wrong). Dots along the top are chains that produce a correct final answer; dots along the bottom are chains with a wrong final answer. The shaded region marks the “mostly correct but wrong” zone: additive reward $\geq 0.80$ yet validity =0=0. Eight chains (16%) fall in this zone. Six of these have every step rated correct (additive reward =1.0=1.0) but still produce a wrong final answer, because the step-level heuristic evaluates numerical consistency of intermediate calculations but cannot catch errors in problem setup or final answer extraction. Panel (b) shows the distribution of additive rewards separately for correct-answer and wrong-answer chains; the dashed line marks the $0.80$ threshold. Both correct-answer and wrong-answer chains concentrate at high additive reward, confirming that step-level scores do not reliably separate valid from invalid reasoning. These chains span 23–48 reasoning steps, so the phenomenon is not an artifact of short chains. An additive objective would assign near-maximal reward to all eight, reinforcing the reasoning patterns that produce wrong answers; the validity indicator correctly identifies them as failures.
[…]
Validation of the Objective Mismatch (Definition 5.1). (a) Additive vs. multiplicative reward for 50 GSM8K chains. The shaded region highlights 8 chains (16%) that are “mostly correct” (high additive reward) but wrong (multiplicative reward =0=0). (b) Distribution of step correctness, showing that wrong-answer chains can have nearly all steps correct.
8 Conclusion
This paper establishes information-theoretic limits on credit assignment in deep sequential processes, revealing a phase transition that constrains both inspection system design and reasoning model training.
The core insight.
The Signal Decay Bound (Theorem 3.9) shows that information about step $t$ decays as $\eta^{H-t}$, defining a critical horizon:
$H_{\mathrm{crit}}=\frac{\ln(n\Delta^{2})}{\ln(1/\eta)}.$
Steps within $H_{\mathrm{crit}}$ of the outcome can be evaluated from endpoint data; steps beyond cannot, regardless of sample size or algorithmic sophistication.
Why this matters for AI reasoning.
The framework explains why outcome-trained value networks collapse to near-constant predictions on long reasoning chains (Kazemnejad et al. 2024): this is statistically optimal behavior when outcome labels carry negligible information about early steps. Token-level dynamics are deterministic, but credit assignment targets semantic content, and the many-to-one abstraction from tokens to meaning is where contraction arises (see Section 14 in the Electronic Companion for a detailed treatment). Process reward models work by providing signal beyond the critical horizon; width-based methods like GRPO (DeepSeek-AI 2024) face diminishing returns because correlation caps effective width at $1/\rho$ (Theorem 4.4).
Beyond AI: unifying inspection design.
The same mathematics applies across operations settings (manufacturing, service operations, supply chains) wherever sequential processes are evaluated at endpoints. The critical horizon formula provides a direct diagnostic for whether intermediate inspection is necessary, and the optimal scheduling results (Section 6) show that uniform spacing is minimax-optimal under homogeneous contraction, with a greedy algorithm handling heterogeneous contraction. Achieving polynomial sample complexity requires inspection density scaling as $\Omega(H/\log H)$ (Corollary 6.16). A separate failure mode arises from objective structure: standard additive reward aggregation optimizes expected correct steps rather than end-to-end validity (Section 5), and the curriculum of Proposition 5.3 bridges this gap by interpolating between learnable and correct objectives as policy quality improves.
Limitations and future directions.
Several limitations merit acknowledgment. The contraction coefficient $\eta$ is typically unknown and must be estimated empirically. The semantic state abstraction is posited conceptually but not constructively defined; developing practical methods to identify semantic states from token sequences remains open. Binary validity is a simplification; real reasoning admits degrees of correctness, partial credit, and recoverable errors.
Extending the framework in three directions would strengthen its practical impact: (1) graded validity models that distinguish fatal errors from recoverable mistakes; (2) empirical estimation procedures for $\eta$ and $\Delta^{2}$ from observational data; and (3) adaptive inspection schemes that learn contraction parameters online while optimizing checkpoint placement.
In summary.
Our results quantify exactly when outcome supervision suffices and when intermediate feedback is provably necessary:
Condition Implication
$H<H_{\mathrm{crit}}$ Outcome supervision suffices
$H>H_{\mathrm{crit}}$ Process supervision required
$\rho>0$ Effective width capped at $1/\rho$
Additive objective Optimizes wrong quantity for validity
Homogeneous contraction Uniform checkpoint spacing is minimax-optimal
By establishing what endpoint data can and cannot reveal, this work provides a common analytical foundation for inspection design in operations and supervision design in AI.
9 Code and Data Disclosure
The code and data to support the numerical experiments in this paper will be made publicly available upon acceptance.