### v2

[…]
credit assignment, critical horizon, deep reasoning, process supervision, signal decay, inspection design
1 Introduction
When a semiconductor wafer fails final testing, which of the hundred preceding fabrication steps caused the defect? When a customer churns after a multi-touchpoint journey, which interaction was responsible? When a multi-tier supply chain delivers a defective product, which supplier introduced the flaw? These are instances of the credit assignment problem: attributing terminal outcomes to intermediate stages in deep sequential processes.
Classical inspection theory (Lindsay and Bishop 1964, Raz 1986, Tapiero 1987) addresses this challenge through intermediate checkpoints, optimizing their placement to balance inspection cost against defect-escape risk. But intermediate inspection is not always an option: a proof step that appears irregular may be valid in context; a component may only be testable once assembled; a customer’s reaction to one touchpoint depends on what follows. In such cases, validity is a property of the whole sequence, and terminal inspection is not merely cheaper but necessary. Moreover, even when terminal inspection is the only option, the classical literature implicitly assumes that defect signals persist through the production chain: with enough terminal observations, any stage’s quality can eventually be inferred. We show this assumption fails for deep processes, where signals decay exponentially with depth.
The challenge of attributing outcomes to early stages under terminal-only observation has gained particular urgency in AI, where large language models generate extended reasoning chains but only the final answer can be verified. Empirically, reinforcement learning (RL) value networks trained on outcome rewards “collapse”: they converge to near-constant predictions for early steps, as if outcomes carry no information about the chain’s beginning (Kazemnejad et al. 2024). Process reward models that provide step-level supervision dramatically improve performance (Lightman et al. 2024), but require expensive human annotation. When is such annotation necessary, and when does outcome supervision suffice?
This paper answers these questions by establishing information-theoretic foundations for credit assignment. Using Strong Data Processing Inequalities (SDPIs), we prove that signals connecting early steps to terminal outcomes decay exponentially with depth. The analysis reveals a sharp threshold: there exists a critical horizon $H_{\mathrm{crit}}$ such that steps within this distance of the outcome can be evaluated from endpoint data, while steps beyond require a number of samples that grows exponentially in the gap HtH-t. This is not an algorithmic limitation but an information-theoretic barrier. The critical horizon scales logarithmically with sample size and signal strength but inversely with a per-stage contraction rate that captures how much each processing step blurs the distinction between good and bad earlier steps. In AI reasoning systems, sampling temperature controls this blurring: higher temperature produces more diverse continuations, making it harder to trace outcomes back to early steps. In manufacturing, the analogous driver is the overlap among downstream quality-grade distributions: variability-reducing steps (polishing, annealing, rework), pooling operations (lot mixing, kitting, assembly), and coarse metrology bins all cause products of different upstream grades to produce similar distributions of grades at the next stage, fading the signal about the original grade. The critical horizon formula thus determines whether intermediate feedback is a cost-saving measure or a feasibility requirement.
A key insight underlying our analysis is that credit assignment targets abstract states, not raw observations. In manufacturing, the observable state at production stage $t$ consists of detailed sensor readings (temperatures, pressures, vibration spectra), but credit assignment asks whether the stage produced output that was within specification, marginal, or defective. Multiple sensor patterns correspond to the same quality grade. In AI reasoning, the observable state is the token sequence generated after $t$ reasoning steps, but credit assignment asks whether the logical content is valid, not which specific words were used. The phrases “Therefore x=5x=5” and “Thus we conclude $x$ equals 55” use different tokens but express the same mathematical content. In both settings, the space of abstract states (quality grades, semantic content) is smaller than the space of raw observations (sensor readings, token sequences).
This many-to-one mapping from observations to abstract states is the mechanism behind the per-stage contraction introduced above: when multiple observable states collapse to the same abstract state, transitions from distinct starting points lead to overlapping distributions over next states, and statistical distinguishability shrinks. This occurs even when raw dynamics are deterministic (as with token generation in language models), because the induced dynamics on the abstract space contract. The implication for learning systems is immediate: when signal decay is severe, outcome distributions conditioned on different early states become nearly identical, and training data contains almost no information about which early states are better. Under the assumptions of our model, the critic collapse phenomenon observed in AI systems (Kazemnejad et al. 2024) is statistically optimal behavior given the available information, and not a bug to be fixed but a symptom of a fundamental limit.
A natural response to signal decay is to generate more observations: multiple parallel trajectories from the same starting point, as in replicated testing or width-based methods like GRPO (Group Relative Policy Optimization; DeepSeek-AI 2024). We show this strategy provides only limited relief: correlation among rollouts caps the effective width at $1/\rho$, where $\rho$ is the pairwise correlation between rollout outcomes, and adding more rollouts yields diminishing returns.
For steps beyond the critical horizon, process supervision becomes necessary, raising a new question: how should step-level feedback be aggregated? Sequential validity requires all steps to be correct, a multiplicative rather than additive criterion. The mismatch is stark: a 100-step process with 99% per-step accuracy has an expected correctness score of 99, yet the probability that all 100 steps are correct is only $0.99^{100}\approx 0.37$. Systems that optimize additive objectives learn to produce chains that are “mostly correct” but nonetheless invalid. We analyze this objective mismatch and propose curriculum strategies that transition from learnable additive objectives to correct multiplicative ones as policy quality improves.
Once the need for process supervision is established, a natural question arises: where should intermediate checkpoints be placed? Each checkpoint resets the signal decay clock, so a step immediately before a checkpoint is easy to evaluate, while a step far from any checkpoint is hard. The hardest step to evaluate under any schedule is the one farthest from its next checkpoint, and the design goal is to make this worst case as mild as possible. We derive inspection schedules that are minimax-optimal in this sense: they minimize the worst-case sample complexity across all steps. When all stages attenuate signal at the same rate, uniform checkpoint spacing is minimax-optimal. When different stages attenuate at different rates, a greedy algorithm that places each successive checkpoint as far as the information budget allows yields optimal non-uniform schedules, automatically concentrating inspections where signal loss is fastest. In either case, achieving polynomial sample complexity requires inspection density that scales nearly linearly with process depth: deeper processes demand proportionally more intermediate verification.
We develop a detailed application to large language model reasoning (Section 18 in the Electronic Companion), where the state abstraction maps token sequences to semantic content and sampling temperature controls the contraction coefficient. We validate the theoretical predictions through numerical experiments on both synthetic Markov chains and LLM reasoning chains on GSM8K (a grade-school math benchmark); full experimental details appear in Appendix 17.
Summary of contributions.
We establish four main results that transform the intuition “credit assignment in long processes is hard” into precise, actionable principles.
  • Signal Decay Bound (Section 3): Information about any given step decays exponentially when propagating to the terminal outcome. The resulting sample complexity bounds are tight, establishing that exponential scaling is fundamental. This yields a closed-form critical horizon formula.
  • Width Limits (Section 4): Parallel rollouts reduce estimator variance, but correlation caps effective width at $1/\rho$, extending the critical horizon only logarithmically.
  • Objective Mismatch (Section 5): Additive reward aggregation optimizes the wrong quantity when sequential validity requires all steps to be correct. We propose curriculum strategies that interpolate between learnable and correct objectives.
  • Optimal Inspection Design (Section 6): Uniform checkpoint spacing is minimax-optimal under homogeneous contraction; a greedy algorithm yields optimal non-uniform schedules under heterogeneous contraction. Polynomial sample complexity requires inspection density scaling as $\Omega(H/\log H)$.
The framework yields concrete guidance for both AI and operations practitioners. For AI systems, it explains why process reward models improve performance (they provide signal beyond the critical horizon), why methods like VinePPO (value-based intermediate-step process supervision; Kazemnejad et al. 2024) succeed (they reduce effective depth below $H_{\mathrm{crit}}$), and why width-based methods face diminishing returns (correlation caps effective width). For operations managers, it provides a direct diagnostic. Consider a 50-stage semiconductor process. With contraction rate $\eta=0.85$, signal strength $\Delta^{2}=0.2$, and 10,000 test wafers, $H_{\mathrm{crit}}\approx 47$ (simplified; $\approx 48$ with $\epsilon=0.1$): at least one intermediate checkpoint is required. With $\eta=0.90$, the critical horizon extends to 72 steps, and terminal testing alone suffices. (In practice, $\eta$ is auditable: estimate adjacent-stage grade transition matrices from production data and compute the overlap-based bounds in Section 3.2.)
The unifying insight is that credit assignment difficulty is a structural barrier, not an algorithmic one. The critical horizon formula answers the diagnostic question (when is intermediate feedback necessary?); the optimal inspection schedules answer the prescriptive question (where to place checkpoints?); and the objective mismatch analysis answers the aggregation question (how to combine step-level feedback?). Together, these results provide a principled foundation for supervision system design in both operations and AI.
1.1 Literature Review
Our results sit at the intersection of inspection design in operations research, outcome-versus-process supervision in deep sequential decision systems, and information-theoretic limits for Markovian signal propagation. On the OR side, we build on classical checkpoint placement models (Lindsay and Bishop 1964, Raz 1986, Tapiero 1987), but focus on a different binding constraint: observability under terminal-only evaluation in deep processes, rather than expected-cost tradeoffs under persistent defect signals. On the AI side, recent work documents that outcome-only supervision can fail on long reasoning chains and that step-level supervision can help (Lightman et al. 2024, Uesato et al. 2022, Kazemnejad et al. 2024); we provide a non-asymptotic, information-theoretic characterization of when outcome supervision becomes insufficient. In standard RL settings with bounded concentrability, outcome and process supervision are polynomially equivalent (Jia et al. 2025); our setting falls outside this regime because concentrability grows exponentially in horizon. Independent concurrent work by chen2025exponential establishes an exponential separation using function approximation complexity, complementing our information-theoretic approach.
More broadly, this work contributes to a growing OR-for-AI agenda that applies operations research methodology to AI system design (Snell et al. 2025, Besta et al. 2025, Elmachtoub and Grigas 2022, Wiberg et al. 2025), here showing that classical tools such as strong data processing inequalities and minimax inspection design yield precise answers to questions the AI community has addressed primarily through empirical heuristics. Technically, the analysis leverages strong data processing inequalities (Polyanskiy and Wu 2017, Raginsky 2016) and Le Cam-style hypothesis testing lower bounds (Le Cam 1973) to establish tight sample-complexity limits and the closed-form critical horizon. A detailed, stream-by-stream literature discussion (inspection design, RL theory, information theory, and OR-for-AI) is provided in the Electronic Companion (Appendix 10).
1.2 Paper Organization
The paper develops a logical progression from information-theoretic limits to optimal inspection design. Section 2 establishes the formal framework: the MDP model, state abstraction, supervision regimes, and the contraction property that drives signal decay. Section 3 proves the Signal Decay Bound, showing that information about step $t$ decays as $\eta^{H-t}$, and derives the critical horizon formula that determines which steps can be evaluated from terminal data.
Two extensions follow. Section 4 analyzes width amplification, showing that correlation among parallel rollouts caps effective width and extends the critical horizon only logarithmically. Section 5 addresses an orthogonal failure mode: standard additive objectives optimize the wrong quantity for sequential validity, and proposes curriculum strategies that interpolate between learnable and correct objectives.
Section 6 synthesizes these results into optimal inspection schedules: uniform spacing under homogeneous contraction and a greedy information-distance algorithm under heterogeneous contraction. Section 7 provides a concise summary of numerical validation; full experimental setups and all plots are deferred to the Electronic Companion (Appendix 17). The Electronic Companion also develops a detailed application to language model reasoning (Section 18).
2 Problem Formulation
We formalize the credit assignment problem in sequential processes, beginning with the operational setting before developing the mathematical framework.
2.1 Multi-Stage Processes with Terminal Inspection
Consider a process where items pass through $H$ stages before final inspection. At each stage, the item’s state changes and defects may be introduced. The final inspection reveals whether the complete item meets specifications but does not identify which stage caused a failure.
As discussed in Section 1, terminal-only inspection arises either because intermediate inspection is costly (semiconductor metrology, supplier auditing) or infeasible (validity depends on context that emerges only downstream). The common structure across manufacturing, service operations, supply chains, and AI reasoning is that terminal outcomes are observable but the intermediate stage responsible for success or failure is not. Without localization, process improvement efforts may target the wrong stages.
Our framework applies when (i) the process evolves through discrete stages with Markovian dynamics on an abstract state space, (ii) terminal inspection reveals aggregate quality, and (iii) intermediate inspection, if feasible, is costly. In AI, outcome supervision refers to training on final output labels, while process supervision provides step-level feedback. We adopt this terminology throughout, noting that the underlying structure is domain-general.
2.2 Mathematical Formalization
We model the sequential process as an episodic finite-horizon Markov decision process (MDP) with terminal reward. Let $\mathcal{S}$ denote a state space, $\mathcal{A}$ an action space, and $H$ the horizon (the number of decision stages).
Definition 2.1 (Finite-Horizon MDP with Terminal Reward)
A finite-horizon MDP with terminal reward consists of:
  • An initial state $S_{0}\in\mathcal{S}$, drawn from a distribution $\mu$ over $\mathcal{S}$ (or fixed at a given state $s_{0}$)
  • Transition dynamics, either deterministic $s_{t+1}=f(s_{t},a_{t})$ or stochastic $s_{t+1}\sim P(\cdot\mid s_{t},a_{t})$
  • A policy $\pi=(\pi_{0},\ldots,\pi_{H-1})$ where each $\pi_{t}:\mathcal{S}\to\Delta(\mathcal{A})$ maps states to distributions over actions
  • A terminal evaluation $R=g(S_{H})\in\{0,1\}$ indicating success or failure
  • Step validity indicators $r_{t}=h_{t}(S_{t},A_{t},S_{t+1})\in\{0,1\}$, treated as exogenous labels available only under process or partial supervision; the functions $h_{t}$ are problem-dependent and need not be known to the learner
A trajectory is the complete sequence $\tau=(S_{0},A_{0},S_{1},\ldots,A_{H-1},S_{H})$. We use uppercase letters for random variables; realized samples are denoted $\tau^{(i)}=(s_{0}^{(i)},a_{0}^{(i)},\ldots,s_{H}^{(i)})$. In sequential validity tasks, terminal success requires every step to be correct: $R=\prod_{t=0}^{H-1}r_{t}$; this structural relationship motivates the multiplicative objective introduced in Section 5.
To ground this abstraction: in manufacturing, the state $S_{t}$ is the workpiece condition at stage $t$, an action $A_{t}$ is the process setting (temperature, pressure, feed rate), and the transition $f$ maps current condition and setting to the next-stage condition. The terminal evaluation $R$ indicates whether the finished product passes final inspection. In AI reasoning, $S_{t}$ is the partial solution generated so far, $A_{t}$ is the next reasoning step, and $R$ indicates whether the final answer is correct.
Remark 2.2 (Deterministic Dynamics with Stochastic Trajectories)
In many sequential processes, the transition function $f$ is deterministic while trajectories are stochastic because actions are sampled from a policy distribution. This distinction matters: information loss arises not from noisy dynamics but from the many-to-one structure of state abstraction (Section 3). In manufacturing, this occurs even under fixed recipes: units sharing the same quality grade may differ in latent characteristics (material lot, tool drift, defect mode), so the abstraction from detailed workpiece state to grade is many-to-one, and the induced grade-to-grade kernels have overlapping rows and $\eta<1$.
The value function captures the probability of eventual success from any intermediate state. Intuitively, it answers the question: “If the process is currently at state $s$ after $t$ steps, what is the chance it will ultimately succeed?”
Definition 2.3 (Value Function)
The value function under policy $\pi$ from state $s$ at step $t$ is
$V^{\pi}_{t}(s):=\mathbb{P}_{\pi}(R=1\mid S_{t}=s)=\mathbb{E}_{\pi}[R\mid S_{t}=s],$}
representing the probability of terminal success when starting from state $s$ at step $t$ and following policy $\pi$.
2.3 Supervision Models
The learner observes $n$ i.i.d. trajectories generated by a fixed behavior policy $\pi$. We define three supervision regimes, distinguished by what feedback accompanies each trajectory.
Definition 2.4 (Outcome Supervision)
Under outcome supervision, the learner observes $n$ independent trajectory-outcome pairs:
$\mathcal{D}_{\mathrm{out}}=\{(\tau^{(i)},R^{(i)})\}_{i=1}^{n}.$}
The full trajectory $\tau^{(i)}$ including all intermediate states is observed, but no step-level correctness labels are provided.
The learner sees what happened at each step but not whether each step was correct. Consequently, the only supervision signal about step validity is through $R$, and any test for whether step $t$ was correct must rely on the induced separation in outcome distributions; this is why the lower bounds in Section 3 are stated in terms of $P^{(0)}_{R}$ versus $P^{(1)}_{R}$.
Definition 2.5 (Process Supervision)
Under process supervision, the learner additionally observes step validity labels:
$\mathcal{D}_{\mathrm{proc}}=\{(\tau^{(i)},r_{0}^{(i)},\ldots,r_{H-1}^{(i)},R^{(i)})\}_{i=1}^{n},$}
where $r_{t}^{(i)}\in\{0,1\}$ indicates whether step $t$ of trajectory $i$ is valid.
Process supervision provides the ground truth about each step. This is the gold standard for credit assignment but is expensive: a human expert must evaluate every step of every trajectory, or a formal verifier must check each logical inference.
Definition 2.6 (Partial Supervision)
Under $k$-step supervision, the learner observes step labels at intervals of $k$ steps:
$\mathcal{D}_{k}=\{(\tau^{(i)},r_{k}^{(i)},r_{2k}^{(i)},\ldots,r_{\lfloor(H-1)/k\rfloor\cdot k}^{(i)},R^{(i)})\}_{i=1}^{n},$}
That is, labels are provided at steps $t=k,2k,\ldots$ (we omit $t=0$ by convention). This interpolates between near-process supervision ($k=1$, all steps except $r_{0}$ labeled) and outcome supervision ($k\geq H$, no step labels). The convention ensures that coarse supervision provides no step-level information.
Under all three regimes, the learner observes the complete state sequence $(s_{0},\ldots,s_{H})$. The difficulty is not hidden states but missing correctness labels: the model’s output is fully readable, but determining whether each step is valid may require human experts or formal verification.
2.4 Information-Theoretic Preliminaries
Our lower bounds rely on tracking how information degrades through sequential processing. The core question is: if two hypotheses (say, “step $t$ was good” versus “step $t$ was bad”) lead to different distributions over trajectories, how many samples suffice to distinguish them?
The answer depends on how different the resulting outcome distributions are: nearly identical distributions require many samples, while well-separated distributions require few. We formalize this using $\chi^{2}$-divergence.
Definition 2.7 ($\chi^{2}$-Divergence)
For distributions $P$ and $Q$ with $P$ absolutely continuous with respect to $Q$:
$\chi^{2}(P\|Q):=\int\frac{(dP-dQ)^{2}}{dQ}=\mathbb{E}_{Q}\left[\left(\frac{dP}{dQ}-1\right)^{2}\right].$}
We use $\chi^{2}$-divergence rather than KL divergence for two reasons. First, it connects directly to hypothesis testing: by Le Cam’s method (Le Cam 1973), distinguishing $P$ from $Q$ with constant probability requires $n=\Omega(1/\chi^{2}(P\|Q))$ samples. Second, it tensorizes cleanly across independent samples: $\chi^{2}(P^{\otimes n}\|Q^{\otimes n})=(1+\chi^{2}(P\|Q))^{n}-1$ (see, e.g., Polyanskiy and Wu 2017). This multiplicative structure translates divergence bounds directly into sample complexity bounds.
The central tool in our analysis is the Strong Data Processing Inequality, which captures a fundamental phenomenon: processing information can only destroy it, never create it. If two distributions are hard to distinguish, passing them through a noisy channel makes them even harder to distinguish.
Definition 2.8 ($\chi^{2}$-Contraction Coefficient)
For a Markov kernel $K:\mathcal{X}\to\Delta(\mathcal{Y})$, the $\chi^{2}$-contraction coefficient is
$\eta_{\chi^{2}}(K):=\sup_{\begin{subarray}{c}P,Q:P\ll Q\\ 0<\chi^{2}(P\|Q)<\infty\end{subarray}}\frac{\chi^{2}(PK\|QK)}{\chi^{2}(P\|Q)},$}
where $PK(y)=\int K(y\mid x)\,dP(x)$ is the pushforward measure.
The contraction coefficient satisfies $\eta_{\chi^{2}}(K)\in[0,1]$, measuring what fraction of distinguishability survives the channel: $\eta=0$ destroys all information; $\eta=1$ is lossless. Throughout, smaller $\eta$ means stronger contraction (faster information loss).
Lemma 2.9 ($\chi^{2}$ Strong Data Processing Inequality (see Raginsky 2016, Polyanskiy and Wu 2017))
Let $K:\mathcal{Y}\to\Delta(\mathcal{Z})$ be a Markov kernel. For any distributions $P_{Y}$ and $Q_{Y}$ on $\mathcal{Y}$ with $P_{Y}\ll Q_{Y}$ and $\chi^{2}(P_{Y}\|Q_{Y})<\infty$, define $P_{Z}=P_{Y}K$ and $Q_{Z}=Q_{Y}K$. Then
$\chi^{2}(P_{Z}\|Q_{Z})\leq\eta_{\chi^{2}}(K)\cdot\chi^{2}(P_{Y}\|Q_{Y}).$}
For credit assignment: if $P_{Y}$ and $Q_{Y}$ represent state distributions under “good step” versus “bad step,” and $K$ represents subsequent reasoning, then outcome distributions are at most $\eta$ times as distinguishable as the original states. When we iterate through multiple steps, information loss compounds.
Corollary 2.10 (Divergence Decay in Markov Chains (cf. Polyanskiy and Wu 2017))
Consider a Markov chain $(S_{0},S_{1},\ldots,S_{H})$ with transition kernels $K_{0},K_{1},\ldots,K_{H-1}$. For any two initial distributions $P_{S_{0}}$ and $Q_{S_{0}}$ with $P_{S_{0}}\ll Q_{S_{0}}$ and $\chi^{2}(P_{S_{0}}\|Q_{S_{0}})<\infty$:
$\chi^{2}(P_{S_{H}}\|Q_{S_{H}})\leq\left(\prod_{t=0}^{H-1}\eta_{\chi^{2}}(K_{t})\right)\chi^{2}(P_{S_{0}}\|Q_{S_{0}}).$}
In particular, if $\eta_{\chi^{2}}(K_{t})\leq\eta<1$ for all $t$, then
$\chi^{2}(P_{S_{H}}\|Q_{S_{H}})\leq\eta^{H}\cdot\chi^{2}(P_{S_{0}}\|Q_{S_{0}}).$}
3 The Signal Decay Bound: Why Early Steps Are Hardest
We now prove that information about step $t$ decays exponentially as it propagates through the remaining HtH-t stages, making credit assignment to early steps fundamentally harder than to late ones. We first introduce a state abstraction under which abstract dynamics are Markov (Section 3.1), show that the many-to-one structure of this abstraction induces contraction (Section 3.2), and then prove the Signal Decay Bound, derive the critical horizon, and establish tightness (Section 3.3). Proofs appear in Appendix 11.
3.1 State Abstraction and Abstract Dynamics
Credit assignment targets abstract content, not raw observations. In many sequential processes, the full observable state contains details (surface representation, sensor noise, formatting choices) that are irrelevant to the outcome. Multiple observable states may map to the same quality grade or semantic content. This many-to-one structure is the source of information loss, and formalizing it is the first step toward quantifying signal decay.
Definition 3.1 (Markov State Abstraction)
A Markov state abstraction is a mapping $\phi:\mathcal{S}\to\mathcal{Z}$ from observable states to an abstract state space $\mathcal{Z}$ satisfying the lumpability condition: for all $t$ and all $s,s^{\prime}\in\mathcal{S}$ with $\phi(s)=\phi(s^{\prime})=z$,
$\mathcal{L}(\phi(S_{t+1})\mid S_{t}=s)=\mathcal{L}(\phi(S_{t+1})\mid S_{t}=s^{\prime}).$}
The abstract state at step $t$ is $Z_{t}=\phi(S_{t})$.
The lumpability condition, due to Kemeny and Snell (1960), ensures that the abstract process $(Z_{0},Z_{1},\ldots,Z_{H})$ is itself Markov: the distribution of the next abstract state depends only on the current abstract state, not on which observable state within the equivalence class was realized. When this condition holds, there exist well-defined abstract transition kernels $K_{t}:\mathcal{Z}\to\Delta(\mathcal{Z})$ with $K_{t}(z^{\prime}\mid z)=\mathbb{P}(\phi(S_{t+1})=z^{\prime}\mid\phi(S_{t})=z)$.
Lumpability is a modeling choice rather than a restrictive hypothesis: one defines $\phi$ so that the abstract dynamics are Markov. If an initial choice of $\phi$ fails the condition, one refines the partition by augmenting $\phi$ with additional context until the condition holds, at the cost of a larger abstract space. The interesting question is not whether lumpability holds (it can always be arranged, at the cost of enlarging the abstract state space) but how much the resulting abstract kernels $K_{t}$ mix: do different abstract states lead to overlapping or distinct distributions over successor states? A coarser abstraction lumps more observable states together, increasing overlap and accelerating information loss. A finer abstraction preserves more distinctions but may require a larger state space. Section 3.2 formalizes this tradeoff. Section 18 in the Electronic Companion develops a concrete instantiation for language model reasoning.
Example 3.2 (State Abstraction in Applications)
  • Manufacturing: The observable state $S_{t}$ consists of sensor readings at production stage $t$ (temperatures, pressures, vibration spectra). The abstract state $Z_{t}=\phi(S_{t})$ captures the quality grade: “within specification,” “marginal,” or “defective.” Multiple sensor patterns map to the same grade.
  • AI reasoning: The observable state $S_{t}$ is the token sequence generated after $t$ reasoning steps. The abstract state $Z_{t}=\phi(S_{t})$ captures semantic content: the mathematical claims established, the proof strategy adopted, or the intermediate values derived. Multiple token sequences expressing the same content map to the same abstract state.
The lumpability condition requires exact agreement of abstract transition probabilities within each equivalence class. In practice, this may hold only approximately: two sensor patterns classified as the same quality grade may lead to slightly different distributions over downstream grades. We return to this point in Section 3.3, after establishing the main bound, where we show that approximate lumpability incurs only a bounded, typically negligible penalty.
3.2 Why Abstract Dynamics Contract
The many-to-one structure of $\phi$ introduced in Section 3.1 causes information loss whenever distinct abstract states lead to overlapping next-state distributions. We now quantify this loss per step using the $\chi^{2}$-contraction coefficient $\eta_{\chi^{2}}(K)\in[0,1]$ (Definition 2.8), which measures what fraction of $\chi^{2}$-divergence survives when distributions pass through a kernel $K$. When different abstract states $z$ and $z^{\prime}$ produce overlapping distributions $K(\cdot\mid z)$ and $K(\cdot\mid z^{\prime})$ over next states, $\eta_{\chi^{2}}(K)<1$ and any statistical difference between two distributions over $Z_{t}$ shrinks after one transition.
To build intuition, suppose we want to determine whether the abstract chain was in state $z$ or state $z^{\prime}$ at step $t$ by observing the state at step $t+1$. If $K(\cdot\mid z)$ and $K(\cdot\mid z^{\prime})$ have disjoint support, observing $Z_{t+1}$ perfectly identifies $Z_{t}$, and no information is lost. But if both distributions assign positive probability to some state $y$, then observing $Z_{t+1}=y$ leaves residual uncertainty about whether $Z_{t}$ was $z$ or $z^{\prime}$. The greater the overlap between $K(\cdot\mid z)$ and $K(\cdot\mid z^{\prime})$, the more information is lost at each step.
Example 3.3 (Contraction Coefficient in Extreme Cases)
  • Deterministic dynamics ($\eta=1$): If each state $z$ transitions to a unique successor, the transition matrix is a permutation and $K(\cdot\mid z)$ and $K(\cdot\mid z^{\prime})$ have disjoint support for $z\neq z^{\prime}$. No information is lost: $\eta_{\chi^{2}}(K)=1$.
  • Uniform mixing ($\eta=0$): If $K(\cdot\mid z)=\nu$ for all $z$ (every state transitions to the same distribution $\nu$), then $Z_{t+1}$ is independent of $Z_{t}$ and all information about the past is destroyed in one step: $\eta_{\chi^{2}}(K)=0$.
  • Intermediate case: Consider a two-state chain with $K(1\mid 1)=1-p$, $K(2\mid 1)=p$, $K(1\mid 2)=p$, $K(2\mid 2)=1-p$. The overlap between rows is $2\min\{p,1-p\}$, and $\eta_{\chi^{2}}(K)=(1-2p)^{2}$. When $p=0$, the chain is deterministic and $\eta=1$; when $p=1/2$, both rows are identical and $\eta=0$.
The following two propositions formalize the relationship between row overlap and $\chi^{2}$-contraction, providing concrete sufficient conditions for $\eta<1$. Both build on classical Markov chain mixing techniques (Doeblin decomposition, Dobrushin’s ergodicity coefficient) combined with the SDPI framework (Lemma 2.9); proofs appear in Appendix 11.
Proposition 3.4 (Contraction from Diversity)
Consider a transition kernel $K$ on an abstract state space with $|\mathcal{Z}|$ states. If $K$ spreads probability such that $K(z^{\prime}\mid z)\geq p_{\min}>0$ for all state pairs $(z,z^{\prime})$, then $\eta_{\chi^{2}}(K)\leq 1-|\mathcal{Z}|\cdot p_{\min}$ (note $|\mathcal{Z}|\cdot p_{\min}\leq 1$ since rows sum to one).
Proposition 3.4 provides a sufficient but strong condition: requiring $K(z^{\prime}\mid z)\geq p_{\min}$ for all state pairs is often unrealistic. The following proposition gives a weaker condition based on the Dobrushin coefficient (dobrushin1956central), requiring only that each pair of rows has some overlap.
Proposition 3.5 (Contraction from Row Overlap (cf. dobrushin1956central, Polyanskiy and Wu 2017))
For a transition kernel $K$ on a finite abstract state space $\mathcal{Z}$, define the Dobrushin coefficient:
$\alpha(K):=\min_{z,z^{\prime}\in\mathcal{Z}}\sum_{y\in\mathcal{Z}}\min\{K(y\mid z),K(y\mid z^{\prime})\}.$
Then $\eta_{\chi^{2}}(K)\leq 1-\alpha(K)$.
The resulting bound on $\eta$ is looser than Proposition 3.4 but applies more broadly, since it requires only pairwise row overlap rather than full minorization. In practice, $\alpha(K)$ can be estimated by sampling continuations from different prefixes and measuring their distributional overlap. In manufacturing, the same bound is computed from existing instrumentation or pilot audits: bin each stage’s output into grades via $\phi$, estimate $K$ from empirical adjacent-stage transition frequencies, and compute $\alpha(K)$ from the resulting matrix, giving $\eta\leq 1-\alpha(K)$ (Proposition 3.5). Example 15.1 in Appendix 15 illustrates this computation for a three-grade quality transition kernel.
Example 3.3, Example 15.1 in Appendix 15, and Propositions 3.43.5 show that contraction ($\eta<1$) arises whenever the rows of the abstract kernel have nontrivial overlap. Contraction fails (i.e., $\eta=1$ remains possible) only when some pairs of rows have disjoint support ($\alpha(K)=0$), as in deterministic dynamics with distinct successors. We now formalize the assumption that contraction is strict.
{assumption}
[Contractive Dynamics] At each step $u$, the abstract transition kernel $K_{u}$ satisfies $\eta_{\chi^{2}}(K_{u})\leq\eta$ for some $\eta<1$.
Propositions 3.4 and 3.5 provide concrete upper bounds on $\eta$ in terms of the abstract kernel’s row overlap. In Section 18 of the Electronic Companion, we establish how sampling temperature governs $\eta$ in language model reasoning, connecting the abstract contraction framework to a concrete and practically important design lever.
3.3 The Signal Decay Bound
We now state our main result for this section: a lower bound on the number of i.i.d. outcome samples required to distinguish competing hypotheses about early steps. The bound formalizes the intuition that information decays exponentially as it propagates through the sequential chain. We formalize the hypotheses, prove the bound, derive the critical horizon, and show tightness.
Consider two hypotheses $H_{0}$ and $H_{1}$ about step $t$. These hypotheses induce different distributions $P^{(0)}_{Z_{t}}$ and $P^{(1)}_{Z_{t}}$ over the abstract state $Z_{t}$, and consequently different distributions over the binary outcome $R\in\{0,1\}$. We assume the hypotheses agree on everything else: the transition kernels $K_{u}$ for $u\geq t$ and the terminal map $g$ are identical under both. The question is: how many outcome observations are needed to reliably distinguish $H_{0}$ from $H_{1}$?
Definition 3.6 (Hypothesis Model: Localized Distributional Difference)
Fix a step $t\in\{1,\ldots,H\}$. A pair of hypotheses $(H_{0},H_{1})$ differs only at step $t$ if the following generative model holds:
  1. 1.
    Common initial dynamics: Under both hypotheses, the chain $(Z_{0},\ldots,Z_{t-1})$ follows the same distribution determined by the initial state and transition kernels $K_{0},\ldots,K_{t-2}$.
  2. 2.
    Localized divergence: The conditional distribution of $Z_{t}$ given $Z_{t-1}$ differs:
    $H_{i}:\quad Z_{t}\mid Z_{t-1}\sim K_{t-1}^{(i)}(\cdot\mid Z_{t-1}),\quad i\in\{0,1\}.$
    The induced marginal distributions satisfy $\chi^{2}(P^{(0)}_{Z_{t}}\|P^{(1)}_{Z_{t}})=\Delta^{2}>0$.
  3. 3.
    Shared suffix (downstream dynamics): For $u\geq t$, the transition kernels $K_{u}$ and terminal evaluation $g$ are identical under both hypotheses.
In applications: $H_{0}$ corresponds to “step $t$ is valid” (the transition $K_{t-1}^{(0)}$ produces correct states), while $H_{1}$ corresponds to “step $t$ contains an error” (the transition $K_{t-1}^{(1)}$ may produce erroneous states). In manufacturing: $H_{0}$ means stage $t$ operated correctly; $H_{1}$ means a defect was introduced at stage $t$. The shared-suffix condition isolates step $t$ as the sole source of difference: downstream stages process whatever they receive in the same way under both hypotheses, so any difference in the outcome distribution is attributable entirely to step $t$. Errors introduced at step $t$ may be corrected or may cascade through subsequent stages, but the stages themselves operate identically. Appendix 15 discusses how $\Delta^{2}$ can be estimated from pilot audits in operations settings. We index step $t$ as the abstract state $Z_{t}$ immediately after the $t$-th transition, so HtH-t counts the number of transitions remaining to reach the outcome.
Definition 3.7 (Total Testing Error and Distinguishability)
Let $H_{0}$ and $H_{1}$ induce distributions $P^{(0)}$ and $P^{(1)}$ over outcomes. The total testing error, defined as the sum of type-I and type-II error probabilities minimized over all tests, is
$\mathrm{err}^{*}(P^{(0)},P^{(1)}):=\inf_{\psi}\left(\mathbb{P}_{P^{(0)}}(\psi=1)+\mathbb{P}_{P^{(1)}}(\psi=0)\right)=1-d_{\mathrm{TV}}(P^{(0)},P^{(1)}),$}
where the infimum is over all measurable tests $\psi:\mathcal{Y}\to\{0,1\}$. For $n$ i.i.d. samples, the $n$-sample total testing error is $\mathrm{err}^{*}_{n}:=1-d_{\mathrm{TV}}(P^{(0)\otimes n},P^{(1)\otimes n})$, where the test acts on $\mathcal{Y}^{n}$. The hypotheses are $(n,\epsilon)$-distinguishable if $\mathrm{err}^{*}_{n}\leq\epsilon$.
Thus $(n,\epsilon)$-distinguishability requires $d_{\mathrm{TV}}(P^{(0)\otimes n},P^{(1)\otimes n})\geq 1-\epsilon$.
Theorem 3.8 (Signal Decay Bound)
Let $H_{0}$ and $H_{1}$ be two hypotheses differing only at step $t$, inducing distributions $P^{(0)}_{Z_{t}}$ and $P^{(1)}_{Z_{t}}$ over the abstract state with $P^{(0)}_{Z_{t}}\ll P^{(1)}_{Z_{t}}$ and $\chi^{2}(P^{(0)}_{Z_{t}}\|P^{(1)}_{Z_{t}})=\Delta^{2}<\infty$. Suppose the Markov state abstraction (Definition 3.1) and Assumption 3.5 hold with contraction coefficient $\eta<1$, and the binary outcome is determined by the terminal abstract state: $R=g(Z_{H})\in\{0,1\}$. Then:
  1. 1.
    Divergence decay: The $\chi^{2}$-divergence at the outcome satisfies
    $\chi^{2}(P^{(0)}_{R}\|P^{(1)}_{R})\leq\eta^{H-t}\cdot\Delta^{2}.$}
  2. 2.
    Sample complexity lower bound: For any target error $\epsilon\in(0,1/2)$, in the regime $\eta^{H-t}\Delta^{2}\leq 1$, any test achieving total testing error (Definition 3.7) at most $\epsilon$ requires
    $n\geq\frac{(1-\epsilon)^{2}}{\eta^{H-t}\cdot\Delta^{2}}$}
    i.i.d. outcome samples. When $\eta^{H-t}\Delta^{2}>1$, this lower bound becomes $O(1)$; in this regime, the outcome distributions are well-separated and $O(1)$ samples suffice to achieve any constant total error rate.
  3. 3.
    Exponential growth: For fixed initial divergence $\Delta^{2}=\Theta(1)$, the sample complexity grows exponentially in distance to the outcome:
    $n=\Omega\left(\left(\frac{1}{\eta}\right)^{H-t}\right).$}
Remark 11.4 in Appendix 11 extends Theorem 3.8 to noisy terminal observations $R\sim G(\cdot\mid Z_{H})$.
Inverting Theorem 3.8 yields a horizon formula: given a fixed sample budget $n$, how far back can one reliably evaluate?
Corollary 3.9 (Critical Horizon)
For initial divergence $\Delta^{2}$, sample budget $n$, and any target error $\epsilon\in(0,1/2)$, a necessary condition for achieving total testing error at most $\epsilon$ is that the step lies within the $\epsilon$-critical horizon:
$H-t\leq H_{\mathrm{crit}}(\epsilon):=\max\!\left\{0,\;\frac{\ln\!\bigl(n\Delta^{2}/(1-\epsilon)^{2}\bigr)}{\ln(1/\eta)}\right\}.$}
Steps more than $H_{\mathrm{crit}}(\epsilon)$ from the outcome require exponentially many samples to evaluate from outcome labels alone. When $n\Delta^{2}<(1-\epsilon)^{2}$, the horizon is zero: even the final step cannot be distinguished at error $\epsilon$. For intuition, $H_{\mathrm{crit}}\approx\ln(n\Delta^{2})/\ln(1/\eta)$ when $\epsilon$ is a fixed constant.
Example 3.10 (Critical Horizon in Practice)
With $\eta=0.9$, $\Delta^{2}=0.1$, and $n=10^{6}$, the critical horizon is $H_{\mathrm{crit}}=\ln(10^{5})/\ln(1/0.9)\approx 110$: one million outcome samples can reliably evaluate steps up to 110 steps before the outcome. With stronger contraction $\eta=0.7$, the same budget yields $H_{\mathrm{crit}}\approx 32$. The formula depends logarithmically on $n$ and $\Delta^{2}$ but much more strongly on $\eta$: doubling $n$ extends the horizon by $\ln 2/\ln(1/\eta)$ steps ($\approx 6.6$ when $\eta=0.9$), while reducing $\eta$ from $0.9$ to $0.7$ cuts it by a factor of three.
The exponential rate in Theorem 3.8 is achievable: the following proposition constructs a Bernoulli endpoint-testing instance in which the likelihood ratio test matches the lower bound up to constants.
Proposition 3.11 (Achievability: Matching Upper Bound)
Consider a stylized two-hypothesis problem where the outcome $R\in\{0,1\}$ is Bernoulli with parameters:
  • Under $H_{0}$: $\mathbb{P}(R=1)=p_{0}$
  • Under $H_{1}$: $\mathbb{P}(R=1)=p_{1}=p_{0}+\delta$ where $\delta=\Theta(\eta^{(H-t)/2}\Delta)$
Assume $p_{0}\in[\epsilon_{0},1-\epsilon_{0}]$ for some constant $\epsilon_{0}>0$, so the outcome is non-degenerate under both hypotheses. Then the $\chi^{2}$-divergence between outcome distributions is $\chi^{2}(P^{(0)}_{R}\|P^{(1)}_{R})=\Theta(\eta^{H-t}\Delta^{2})$, and the standard likelihood ratio test achieves total testing error $\epsilon$ with sample complexity $n=O(1/(\eta^{H-t}\Delta^{2}))$.
Together, Theorem 3.8 and Proposition 3.11 show that the exponential rate $(1/\eta)^{H-t}$ is both necessary and sufficient for the matched endpoint-testing problem: the scaling is tight up to constants. Example 11.8 in the appendix constructs a full Markov-chain instance achieving this rate with equality, confirming that the barrier is a property of the information structure, not of our proof technique.
The preceding results assume exact lumpability (Definition 3.1), but in practice this may hold only approximately. Proposition 11.9 in Appendix 11 shows that if each step introduces at most $\delta$ total-variation error relative to the lumpable idealization, the total-variation distance between outcome distributions acquires an additive penalty of $2(H-t)\delta$ on top of the signal-decay term $\sqrt{\eta^{H-t}\Delta^{2}/2}$. For a 50-stage process with per-step discrepancy $\delta=0.001$, the accumulated penalty is at most $2\times 50\times 0.001=0.1$ in total variation distance. Approximate lumpability becomes a concern only for very deep processes or coarse abstractions, in which case the partition should be refined.
Theorem 3.8 tells us how many samples are needed for reliable testing, and Corollary 3.9 inverts this into a horizon formula. But neither quantifies how bad the error is when the sample budget falls short. The following corollary fills this gap: it gives an explicit lower bound on the minimax testing error, showing that when the available samples $n$ fall below the threshold $1/(\eta^{H-t}\Delta^{2})$ required by Theorem 3.8, the error remains close to $1/2$; no test can do meaningfully better than random guessing.
Corollary 3.12 (Statistical Limits of Value Estimation)
Fix $\epsilon\in(0,1/2)$ and let $R^{1:n}$ denote $n$ i.i.d. outcome samples. Under the Markov state abstraction (Definition 3.1) and Assumption 3.5, the minimax error for testing $H_{0}$ vs. $H_{1}$ based only on $R^{1:n}$ satisfies
$\inf_{\psi}\max_{i\in\{0,1\}}\mathbb{P}_{i}(\psi(R^{1:n})\neq i)\;\geq\;\frac{1}{2}\left(1-\sqrt{\frac{(1+\eta^{H-t}\Delta^{2})^{n}-1}{2}}\right).$
In particular, if
$n\;\leq\;\frac{\ln\!\left(1+2(1-2\epsilon)^{2}\right)}{\eta^{H-t}\Delta^{2}},$
then every test $\psi$ has minimax error at least $\epsilon$.
Corollary 3.12 uses the per-hypothesis error convention $\max_{i}\mathbb{P}_{i}(\psi\neq i)$, standard for Le Cam bounds, whereas Definition 3.7 uses the total error $\mathbb{P}_{0}(\psi=1)+\mathbb{P}_{1}(\psi=0)$; the two are related by a factor of two and yield equivalent scaling in $\eta^{H-t}\Delta^{2}$. When $n\eta^{H-t}\Delta^{2}$ is small, $(1+\eta^{H-t}\Delta^{2})^{n}\approx 1+n\eta^{H-t}\Delta^{2}$, so the minimax error is close to $1/2$: the best possible test is barely better than a coin flip. For reinforcement learning, this means that critics trained from outcome labels cannot reliably evaluate steps beyond the critical horizon.
4 Limits of Parallel Sampling
The Signal Decay Bound reveals that outcome signals weaken exponentially with depth. A natural response is to generate more samples: $W$ independent observations from the same starting condition, as in replicated testing or width-based methods like GRPO (DeepSeek-AI 2024). The intuition is that even if individual observations carry weak signal, averaging enough of them should recover the truth. This approach might seem to extract more signal from outcome data, despite Section 3 showing fundamental limits.
But can width truly overcome the exponential barrier? This section shows that it cannot: width reduces noise around an estimate but cannot amplify a signal that has already decayed. The statistical tools, variance reduction under averaging and design effects under equicorrelation (kish1965survey), are classical, but their interaction with the signal decay structure of Section 3 yields new limits on the reach of parallel sampling. We proceed in two steps: derive variance reduction bounds under independence (Section 4.1) and show that correlation limits effective width (Section 4.2). Proofs of all results in this section appear in Appendix 12.
4.1 Multi-Rollout Value Estimation
Consider a learner who wishes to estimate the value $V^{\pi}(s):=\mathbb{P}_{\pi}(R=1\mid S_{t}=s)$, the probability of eventual success from state $s$ under the current policy $\pi$. (When conditioning on abstract state, replace $S_{t}$ with $Z_{t}$.) Rather than observing a single trajectory, the learner generates $W$ independent continuations from $s$ and observes their outcomes.
Throughout this section, we analyze the conditional distribution given a realized $s$; unconditional properties follow by averaging over the distribution of $s$.
Definition 4.1 (Multi-Rollout Estimator)
Given a state $s$ and $W$ rollout continuations $\tau_{1},\ldots,\tau_{W}$ from $s$ under policy $\pi$, the multi-rollout value estimator is:
$\hat{V}_{W}(s):=\frac{1}{W}\sum_{j=1}^{W}R_{j},$}
where $R_{j}\in\{0,1\}$ is the binary outcome of rollout $\tau_{j}$.
The estimator is simply the empirical success rate across rollouts. Its statistical properties depend on the independence structure.
{assumption}
[Conditional Independence] Given the conditioning variable $s$ and a fixed policy $\pi$, the rollouts $\tau_{1},\ldots,\tau_{W}$ are conditionally independent:
$P(\tau_{1},\ldots,\tau_{W}\mid s,\pi)=\prod_{j=1}^{W}P(\tau_{j}\mid s,\pi).$}
This assumption holds when each rollout is generated by independent sampling from the policy starting at $s$. It fails when rollouts share randomness, when the policy adapts based on previous rollouts, or when systematic biases cause rollouts to follow similar paths.
Proposition 4.2 (Variance Reduction from Width)
Under Assumption 4.1, the multi-rollout estimator satisfies:
  1. 1.
    Unbiasedness: $\mathbb{E}[\hat{V}_{W}(s)\mid s]=V^{\pi}(s)$.
  2. 2.
    Variance reduction: $\mathrm{Var}(\hat{V}_{W}(s)\mid s)=\dfrac{V^{\pi}(s)(1-V^{\pi}(s))}{W}$.
  3. 3.
    Concentration (Hoeffding): With probability at least $1-\delta$: $|\hat{V}_{W}(s)-V^{\pi}(s)|\leq\sqrt{\frac{\ln(2/\delta)}{2W}}$
The variance in Part 2 is largest when $V^{\pi}(s)=1/2$ (maximum uncertainty) and zero when outcomes are deterministic. Dividing by $W$ confirms the standard result: to halve the standard deviation, quadruple $W$. Part 3 follows from Hoeffding’s inequality.
The preceding analysis assumes independent rollouts. In practice, rollouts from the same starting point are correlated, which further limits the benefit of width.
4.2 Correlation Limits Effective Width
To quantify this limitation, we must account for correlation among rollouts from the same prefix. Consider collecting data from $n$ independent starting points, generating $W$ rollouts from each. Rollouts from different starting points are independent, and their outcomes are uncorrelated. However, rollouts from the same starting point share the same prefix and tend to make similar errors, inducing positive correlation among their outcomes. Correlation arises from shared prefix errors (all continuations inherit a latent early defect), mode collapse (low-entropy policies generate near-identical continuations, driving $\rho\to 1$), and systematic biases (learned error patterns activated by specific inputs). When $W=1$ (one rollout per starting point), we obtain $n$ i.i.d. outcome observations. When $W>1$, this within-starting-point correlation reduces the effective sample size, as the following theorem quantifies.
Theorem 4.3 (Effective Width Under Correlation (cf. kish1965survey))
Consider $W$ rollouts from a single starting point $s$. Assume the outcomes satisfy equicorrelation: $\mathrm{Corr}(R_{i},R_{j}\mid s)=\rho$ for all $i\neq j$, where $\rho\in[0,1)$ is a constant. (More generally, let $\rho$ be an upper bound on pairwise correlations.) The variance of the multi-rollout estimator $\hat{V}_{W}(s)=\frac{1}{W}\sum_{j=1}^{W}R_{j}$ satisfies:
$\mathrm{Var}(\hat{V}_{W}(s)\mid s)=\frac{\sigma^{2}(s)}{W}\cdot(1+(W-1)\rho),$}
where $\sigma^{2}(s)=V^{\pi}(s)(1-V^{\pi}(s))$. Equivalently, the correlated estimator has the same variance as an estimator using $W_{\mathrm{eff}}$ independent rollouts, where the effective width is:
$W_{\mathrm{eff}}:=\frac{W}{1+(W-1)\rho}.$}
When $\rho=0$ (perfect independence), $W_{\mathrm{eff}}=W$ and no degradation occurs. When $\rho>0$, the effective width saturates as $W\to\infty$: $W_{\mathrm{eff}}\to 1/\rho$.
Example 4.4 (Impact of Correlation)
Suppose rollouts have correlation $\rho=0.2$ (modest correlation from shared reasoning patterns). The effective width for various nominal widths:
$W$ 10 50 100 500 $\infty$
$W_{\mathrm{eff}}$ 3.6 4.6 4.8 5.0 5.0
Even with 500 rollouts, the effective width is only 5. Beyond roughly $1/\rho=5$ rollouts, additional sampling provides negligible benefit.
The equicorrelation variance formula in Theorem 4.3 is a classical design-effect result in survey sampling (kish1965survey); what is new is its implication for the credit assignment horizon, which we develop next. While the $W$ rollouts within each starting point are correlated, the $n$ starting points are independent. Define the effective sample size as $N_{\mathrm{eff}}:=n\cdot W_{\mathrm{eff}}$. Combined with the Signal Decay Bound (Theorem 3.8), distinguishing hypotheses about step $t$ requires
$N_{\mathrm{eff}}\;\gtrsim\;\frac{1}{\eta^{H-t}\Delta^{2}}.$}
The following corollary formalizes the resulting critical horizon.
Corollary 4.5 (Critical Horizon with Width)
Let $n$ be the number of independent starting points and $W_{\mathrm{eff}}$ the effective width per starting point. For fixed $\epsilon$, the critical horizon extends to
$H_{\mathrm{crit}}=\frac{\ln(n\cdot W_{\mathrm{eff}}\cdot\Delta^{2})}{\ln(1/\eta)},$
where the $O(\ln(1/(1-\epsilon)))$ correction from Corollary 3.9 is suppressed.
Width enters the critical horizon only logarithmically: doubling $W_{\mathrm{eff}}$ adds merely $\ln(2)/\ln(1/\eta)\approx 6.6$ steps when $\eta=0.9$. And correlation caps the benefit: since $W_{\mathrm{eff}}\leq 1/\rho$, the effective sample size cannot exceed $n/\rho$ regardless of nominal width. The exponential dependence on depth cannot be escaped by increasing $W$.
In sum, width helps only when the step lies within the critical horizon and rollouts are approximately independent; beyond either limit, additional rollouts yield diminishing returns. For deep processes where $H>H_{\mathrm{crit}}$, width cannot substitute for step-level supervision; Appendix 16 provides a decision rule: width alone is provably insufficient whenever $H>\ln(n\Delta^{2}/\rho)/\ln(1/\eta)$, and in practice one inspection dominates doubling $W$ whenever the maximal gap exceeds $H_{\mathrm{crit}}(\epsilon)$. This raises the question of how step-level supervision should be aggregated.
5 From Outcome Limits to Process Supervision
The preceding sections established that outcome-based credit assignment faces an exponential barrier (Theorem 3.8) that width can only logarithmically mitigate (Section 4). For deep processes where $H>H_{\mathrm{crit}}$, this leaves only one path forward: step-level supervision. Process reward models (Lightman et al. 2024), intermediate checkpoints in manufacturing, and stage-gate reviews in project management all implement this principle. Step-level labels are not merely convenient; they are necessary.
But step-level supervision introduces its own challenge: how should step-level feedback be aggregated into a training or evaluation objective? Standard reinforcement learning sums rewards across steps (Sutton et al. 2000), but for sequential validity, where a single error invalidates the entire chain, this additive aggregation optimizes the wrong quantity. Proofs of all results in this section appear in Appendix 13.
5.1 The Objective Mismatch
Given that step-level supervision is necessary for deep processes, how should it be used? When training a system with step-level labels $r_{t}\in\{0,1\}$ indicating whether each step is valid, one can maximize either the expected number of correct steps or the probability that all steps are correct. These correspond, respectively, to additive and multiplicative aggregation of step-level feedback.
Definition 5.1 (Step-Quality and Validity Objectives)
For a policy $\pi$ generating trajectories of length $H$ from initial state $S_{0}$, with step validity indicators $r_{0},\ldots,r_{H-1}$:
  1. 1.
    The step-quality objective is $J_{\mathrm{add}}:=\mathbb{E}_{\pi}\left[\sum_{t=0}^{H-1}r_{t}\,\big|\,S_{0}\right]$, the expected number of correct steps.
  2. 2.
    The validity objective is $J_{\mathrm{mult}}:=\mathbb{E}_{\pi}\left[\prod_{t=0}^{H-1}r_{t}\,\big|\,S_{0}\right]$, the probability that all steps are correct.
Multiplicative objectives appear in risk-sensitive control (Howard and Matheson 1972) and stochastic reachability (Abate et al. 2008). Our contribution is identifying that sequential validity, where a single failure invalidates the entire chain, naturally requires this structure, and analyzing the optimization challenges it creates. This applies to logical reasoning, manufacturing processes where any defective stage produces a defective product, and service operations where any failed touchpoint loses the customer.
Under independent steps with common success probability $p$, we have $J_{\mathrm{add}}=Hp$ while $J_{\mathrm{mult}}=p^{H}$. Even at 99% per-step accuracy over 100 steps, $J_{\mathrm{add}}=99$ while $J_{\mathrm{mult}}=0.37$: systems optimizing additive objectives learn to produce sequences that are mostly correct but nonetheless invalid.
5.2 Gradient Decay in the Multiplicative Objective
The preceding analysis suggests using the multiplicative objective directly. However, this approach encounters a fundamental optimization barrier: the multiplicative objective exhibits vanishing gradients for long chains with imperfect per-step accuracy.
Under the independence model, differentiating $J_{\mathrm{add}}=Hp$ and $J_{\mathrm{mult}}=p^{H}$ with respect to policy parameters $\theta$ gives $\nabla_{\theta}J_{\mathrm{add}}=H\nabla_{\theta}p$ and $\nabla_{\theta}J_{\mathrm{mult}}=Hp^{H-1}\nabla_{\theta}p$. The multiplicative gradient contains an additional factor $p^{H-1}$ that attenuates the learning signal. For $p=0.95$ over a 100-step chain, this factor equals $0.95^{99}\approx 0.006$, rendering the validity objective nearly unlearnable. The barrier becomes severe precisely when it matters most: for longer chains and policies that have not yet achieved high per-step accuracy.
This gradient decay is conceptually distinct from the signal decay of Section 3. Signal decay concerns information flow: outcome data becomes uninformative about early steps. Gradient decay concerns objective structure: the multiplicative objective attenuates learning signal regardless of information availability. A complete training framework must address both barriers. We propose a curriculum approach: begin with the learnable additive objective, then transition to the correct multiplicative objective as policy quality improves (Bengio et al. 2009).
Definition 5.2 (Interpolated Objective)
For $\lambda\in[0,1]$, the interpolated objective is:
$J_{\lambda}(\theta):=(1-\lambda)\cdot J_{\mathrm{add}}(\theta)+\lambda\cdot J_{\mathrm{mult}}(\theta).$}
At $\lambda=0$, the objective is learnable but incorrect; at $\lambda=1$, correct but potentially unlearnable. The following result characterizes how $\lambda$ should evolve with policy quality.
Proposition 5.3 (Gradient-Preserving Curriculum)
Assume $\nabla_{\theta}J_{\mathrm{add}}$ and $\nabla_{\theta}J_{\mathrm{mult}}$ have non-negative inner product. To maintain $\|\nabla_{\theta}J_{\lambda}\|\geq c\cdot\|\nabla_{\theta}J_{\mathrm{add}}\|$ for some $c\in(0,1)$, it suffices to set $\lambda\leq 1-c$.
The proposition establishes that the interpolated objective preserves gradient signal when $\lambda$ is sufficiently small. In the independence model where $J_{\mathrm{add}}=Hp$ and $J_{\mathrm{mult}}=p^{H}$, the gradients satisfy $\|\nabla_{\theta}J_{\mathrm{mult}}\|=p^{H-1}\|\nabla_{\theta}J_{\mathrm{add}}\|$. As the policy improves and per-step success probability $p$ approaches $1$, this attenuation factor approaches $1$, allowing $\lambda$ to increase toward $1$ without sacrificing gradient magnitude. This yields a curriculum: begin training with $\lambda$ near 0 to ensure learnability, then gradually increase $\lambda$ toward $1$ as the policy improves, transitioning from a learnable but misspecified objective to one that is correctly specified.
Having established why step-level supervision is necessary and how to aggregate it, the next section addresses where: given an inspection budget, which steps should be observed?
6 Optimal Intermediate Inspection Design
We now convert the information-theoretic limits into optimal intermediate inspection schedules. The organizing principle is that inspections shorten the information path that credit assignment must traverse: instead of propagating from step $t$ to the terminal outcome at time $H$, it suffices to propagate from $t$ to the next checkpoint.
The inspection design problem has a long history in OR under different guises. In manufacturing, it is the placement of quality checkpoints along a production line (Lindsay and Bishop 1964, Raz 1986). In sequential testing, it is related to adaptive group testing (Du and Hwang 2000). Our contribution is to solve inspection placement under the specific signal-decay structure implied by the Strong Data Processing Inequality, yielding closed-form schedules for both homogeneous and heterogeneous contraction.
We proceed in three steps. First, we formalize the inspection feedback model and extend the signal-decay bound to an arbitrary inspection time $u$. Second, under homogeneous contraction we show that worst-case sample complexity is governed by a single scalar, the largest distance from any step to its next inspection, and we solve the resulting placement problem (uniform spacing). Third, under heterogeneous contraction we show that feasibility depends on cumulative information loss along each segment, and we solve the minimum-inspection design problem via a greedy information-distance schedule. Cost and sample tradeoffs are then handled in Section 6.5. Proofs of all results in this section appear in Appendix 14.
6.1 Inspection Schedules and Feedback Model
Recall the abstract Markov chain $(Z_{0},\dots,Z_{H})$ induced by the abstraction $\phi$ and policy $\pi$. We treat the terminal outcome $R=g(Z_{H})\in\{0,1\}$ as a terminal inspection at time $H$.
Definition 6.1 (Inspection Schedule)
An inspection schedule is a strictly increasing set of inspection times
$\mathcal{T}=\{t_{1}<t_{2}<\cdots<t_{m}\}\subseteq\{1,2,\dots,H-1\}.$
We use the augmented sequence $0=:t_{0}<t_{1}<\cdots<t_{m}<t_{m+1}:=H$. The number of paid intermediate inspections is $m$.
Definition 6.2 (Inspection Observations)
At each inspection time $u\in\mathcal{T}\cup\{H\}$, the learner observes an inspection output $Y_{u}:=g_{u}(Z_{u})\in\mathcal{Y}_{u}$, where $g_{u}$ is an external evaluator (human, verifier, metrology device) applied to the abstract state. The terminal outcome is the special case $Y_{H}\equiv R=g(Z_{H})$.
The model allows many inspection types: prefix-validity checks, milestone verification, partial correctness scores, or any measurement that is a function of the abstract state. Remark 14.2 in Appendix 14 extends the model to noisy inspections with false positives and negatives, adding an observation-channel penalty to the segment attenuation budget. The results below are information-theoretic and do not assume any particular learning algorithm.
6.2 Signal Decay to Intermediate Inspections
The design lever is that intermediate checkpoints reduce the effective distance over which information must propagate. If step $t$ is followed by an inspection at time $u<H$, then credit assignment for step $t$ need only overcome utu-t steps of attenuation rather than HtH-t. We formalize this for inhomogeneous contraction (step-dependent mixing rates); the homogeneous case follows in Section 6.3.
{assumption}
[Inhomogeneous Contractive Dynamics] Let $K_{t}$ denote the abstract transition kernel from $Z_{t}$ to $Z_{t+1}$, with $\chi^{2}$-contraction coefficient $\eta_{\chi^{2}}(K_{t})\leq\eta_{t}\in(0,1]$ for $t=0,1,\ldots,H-1$. Define the cumulative attenuation from step $t$ to inspection time $u>t$ as
$\mathrm{Attn}(t\to u):=\prod_{j=t}^{u-1}\eta_{j}.$
In the homogeneous case $\eta_{t}\equiv\eta$, this simplifies to $\mathrm{Attn}(t\to u)=\eta^{u-t}$.
Proposition 6.3 (Signal Decay to an Intermediate Inspection)
Consider two hypotheses $H_{0}$ and $H_{1}$ differing only at step $t$, inducing abstract state distributions with $\chi^{2}(P^{(0)}_{Z_{t}}\|P^{(1)}_{Z_{t}})=\Delta^{2}$. Under Assumption 6.2, the $\chi^{2}$-divergence at any downstream inspection $Y_{u}=g_{u}(Z_{u})$ satisfies
$\chi^{2}(P^{(0)}_{Y_{u}}\|P^{(1)}_{Y_{u}})\leq\mathrm{Attn}(t\to u)\cdot\Delta^{2}.$
Proposition 6.3 implies that an inspection at time $u$ replaces the terminal attenuation factor $\mathrm{Attn}(t\to H)$ by the shorter-horizon factor $\mathrm{Attn}(t\to u)$. The following corollary makes the value of intermediate inspection precise: placing a checkpoint at time $u$ reduces the attenuation exponent from HtH-t to utu-t. The remaining question is how to place inspection times so that no step is too far from its next checkpoint.
Corollary 6.4 (Sample Complexity with Intermediate Inspection)
Under the conditions of Proposition 6.3, any test distinguishing $H_{0}$ from $H_{1}$ at error rate $\epsilon$ requires
$n\geq\frac{(1-\epsilon)^{2}}{\mathrm{Attn}(t\to u)\cdot\Delta^{2}}.$
In the homogeneous case, $n\geq(1-\epsilon)^{2}/(\eta^{u-t}\Delta^{2})$.
6.3 Homogeneous Contraction: Minimax-Optimal Placement
Under homogeneous contraction, the sample complexity penalty depends only on the distance from a step to its next inspection (Corollary 6.4). This turns inspection design into a one-dimensional placement problem on $\{0,1,\ldots,H\}$ with a single governing design statistic.
The development in this subsection follows a short chain of implications. We first define the distance of each step to its next checkpoint and the induced maximal gap $L(\mathcal{T})$. We then show that any schedule must incur a worst-case sample complexity that depends exponentially on $L(\mathcal{T})$. Finally, we solve the placement problem by minimizing $L(\mathcal{T})$ over schedules with a fixed number of inspections $m$ and translate the result into a minimum number of inspections required for feasibility.
Definition 6.5 (Downstream Distance and Maximal Gap)
For a schedule $\mathcal{T}=\{t_{1}<\cdots<t_{m}\}$ with augmented times $t_{0}=0$ and $t_{m+1}=H$, the downstream distance of step $t$ is the distance to its next inspection:
$d_{\mathcal{T}}(t):=\min\{u-t:u\in\mathcal{T}\cup\{H\},\ u>t\}.$
The maximal gap is $L(\mathcal{T}):=\max_{t\in\{0,\ldots,H-1\}}d_{\mathcal{T}}(t)$, equivalently $L(\mathcal{T})=\max_{i=0,\ldots,m}(t_{i+1}-t_{i})$.
The maximal gap determines worst-case performance: the step at the beginning of the longest segment between inspections faces the most attenuation before any observation. For a fixed number of intermediate inspections $m$, the minimax placement problem is therefore
$\min_{|\mathcal{T}|=m}L(\mathcal{T}),$
because $L(\mathcal{T})$ controls the worst-case attenuation exponent (and hence the worst-case sample complexity) across all steps. Operationally, $L(\mathcal{T})$ is the length of the longest uninspected “run” of steps; the minimax goal is to prevent any stage from being too far from a checkpoint.
Proposition 6.6 (Worst-Case Sample Complexity)
Under homogeneous contraction $\eta_{t}\equiv\eta<1$, any inspection schedule $\mathcal{T}$ admits a step $t^{\star}$ such that distinguishing hypotheses about $t^{\star}$ requires
$n\geq\frac{(1-\epsilon)^{2}}{\eta^{L(\mathcal{T})}\Delta^{2}}$
samples.
Proposition 6.6 identifies the correct minimax objective: reduce the largest gap to reduce worst-case sample complexity. The remaining question is purely combinatorial: how should $m$ inspections partition a horizon of length $H$ to minimize the length of the longest segment?
Theorem 6.7 (Minimax-Optimal Inspection Placement)
Under homogeneous contraction $\eta_{t}\equiv\eta<1$, the minimal achievable maximal gap with $m$ intermediate inspections is
$\min_{|\mathcal{T}|=m}L(\mathcal{T})=\Big\lceil\frac{H}{m+1}\Big\rceil.$
This minimum is achieved by near-uniform spacing: placing inspections at $t_{i}=\lfloor iH/(m+1)\rfloor$ for $i=1,\ldots,m$.
While uniform spacing is a classical heuristic for worst-case gap objectives, our contribution is identifying the correct design objective, namely the gap in SDPI attenuation units, and linking it to sample complexity via the Signal Decay Bound. The implication is that the detailed location of checkpoints is secondary; what matters is controlling the longest uninspected run. Combining Proposition 6.6 and Theorem 6.7 yields an explicit feasibility criterion.
Corollary 6.8 (Sample Complexity Under $m$ Inspections)
With $m$ minimax-optimal inspections, the worst-case sample complexity for credit assignment is
$n\geq\frac{(1-\epsilon)^{2}}{\eta^{\lceil H/(m+1)\rceil}\Delta^{2}}.$
Corollary 6.9 (Minimum Inspections for Feasibility)
For all steps to be $(n,\epsilon)$-distinguishable, the number of inspections must satisfy
$m\geq\Big\lceil\frac{H}{H_{\mathrm{crit}}(\epsilon)}\Big\rceil-1,$
where $H_{\mathrm{crit}}(\epsilon)=\bigl(\ln(n\Delta^{2})-2\ln(1-\epsilon)\bigr)/\ln(1/\eta)$ is the $\epsilon$-critical horizon. This is a necessary condition; the sufficient number may be one higher due to integer rounding when $H/(m+1)$ is not an integer.
Example 15.2 in Appendix 15 applies Corollary 6.9 to a 50-stage semiconductor fabrication process, showing that a single midpoint inspection suffices to bring all stages within the critical horizon.
6.4 Heterogeneous Contraction: Minimum Inspections via Information Distance
Uniform spacing is optimal when all steps contract equally, but many processes exhibit depth-dependent contraction. In such settings, the design objective shifts from “minimize the largest segment length” to a feasibility-driven problem: choose as few inspections as possible so that every segment has cumulative information loss at most the threshold implied by the sample budget.
Definition 6.10 (Information Distance)
Under Assumption 6.2, the per-step information distance is $w_{t}:=\ln(1/\eta_{t})\geq 0$. For an interval $[a,b)$, the cumulative information distance is
$W([a,b)):=\sum_{t=a}^{b-1}w_{t}=\ln(1/\mathrm{Attn}(a\to b)).$
Steps with strong contraction (small $\eta_{t}$) contribute large $w_{t}$; steps with weak contraction contribute little.
Because $\mathrm{Attn}(t\to u)=\exp(-W([t,u)))$, Corollary 6.4 yields a segment-level feasibility threshold.
Lemma 6.11 (Per-Segment Feasibility)
Fix consecutive inspection times $t_{i}<t_{i+1}$. If
$W([t_{i},t_{i+1}))\leq\Gamma(n,\Delta^{2},\epsilon):=\ln(n\Delta^{2})-2\ln(1-\epsilon),$
then every step $t\in\{t_{i},\ldots,t_{i+1}-1\}$ is $(n,\epsilon)$-distinguishable using the inspection at $t_{i+1}$. Conversely, if step $t_{i}$ is $(n,\epsilon)$-distinguishable using the inspection at $t_{i+1}$, then the displayed inequality must hold.
The heterogeneous inspection design problem is to partition the horizon into segments satisfying the Lemma 6.11 constraint while using as few paid inspections as possible. A greedy construction is optimal.
Theorem 6.12 (Greedy Inspection is Optimal)
Among all schedules satisfying $W([t_{i},t_{i+1}))\leq\Gamma$ for each segment, the following greedy construction minimizes the number of inspections: starting at $t_{0}=0$, place each inspection $t_{i+1}$ at the largest index such that $W([t_{i},t_{i+1}))\leq\Gamma$.
Greedy segmentation is standard for interval-packing problems; what the SDPI analysis adds is the metric of cumulative information distance $W([t_{i},t_{i+1}))=\sum_{t}\ln(1/\eta_{t})$ rather than step count, under which feasibility is defined, so inspections concentrate automatically where contraction is strongest. When $\eta_{t}\equiv\eta$, we have $w_{t}\equiv\ln(1/\eta)$, so greedy produces segments of length $\lfloor\Gamma/\ln(1/\eta)\rfloor$, recovering uniform spacing (Theorem 6.7). When contraction varies, greedy produces non-uniform placement: segments spanning high-contraction steps are short; segments spanning low-contraction steps are long. Example 15.3 in Appendix 15 applies Theorem 6.12 to a 50-touchpoint customer journey with depth-dependent contraction, showing that the greedy algorithm places its single inspection in the high-contraction early stages.
6.5 Joint Sample-Inspection Budgeting
Inspection frequency trades off per-trajectory cost against the exponential sample savings from shortening information paths. Let $c_{\mathrm{out}}$ denote the per-trajectory cost of terminal evaluation and $c_{\mathrm{insp}}$ the per-trajectory cost per intermediate inspection. Collecting $n$ trajectories with $m$ inspections each has total budget $B=n(c_{\mathrm{out}}+mc_{\mathrm{insp}})$.
Proposition 6.13 (Budget Lower Bound)
Under homogeneous contraction, the minimum budget required for worst-case $(n,\epsilon)$-distinguishability with $m$ inspections is
$B\geq(c_{\mathrm{out}}+m\cdot c_{\mathrm{insp}})\cdot\frac{(1-\epsilon)^{2}}{\eta^{\lceil H/(m+1)\rceil}\Delta^{2}}.$
Increasing $m$ reduces the sample requirement exponentially but raises per-trajectory cost linearly. A practical rule is to choose the smallest $m$ that makes the uniform segment length $\lceil H/(m+1)\rceil$ fall within $H_{\mathrm{crit}}(\epsilon)$ (Corollary 6.9).
Corollary 6.14 (Inspection Density for Polynomial Complexity)
To achieve polynomial sample complexity $n=O(H^{p})$ under homogeneous contraction, the number of inspections must scale as $m=\Omega(H/\log H)$.
Corollary 6.14 implies that inspection effort must grow nearly linearly with process depth to avoid exponential sample requirements. An implementable inspection design procedure and detailed cost–sample tradeoff analysis are provided in the Electronic Companion (Appendix 16).
7 Numerical Validation
We validate the paper’s five core predictions through synthetic Markov chains with controlled parameters and, where applicable, LLM reasoning chains on GSM8K (a grade-school math benchmark). Across all experiments, observed behavior closely matches the theoretical bounds, confirming that the closed-form results provide reliable quantitative guidance in practice. Table 7 summarizes the headline findings; full experimental setups, parameter configurations, and all figures are provided in the Electronic Companion (Appendix 17).
[…]
Headline numerical findings. Full setups and plots in Appendix 17. Prediction Test bed Key finding Signal decay (Thm. 3.8) 10-state synthetic chains, $\eta\in\{0.7,0.8,0.9,0.95\}$ Measured $\chi^{2}$-divergences match $\eta^{H-t}\Delta^{2}$ exactly ($R^{2}>0.999$). Effective width (Thm. 4.3) Synthetic ($\rho=0.15$); LLM rollouts on GSM8K ($\rho\approx 0.49$) $W_{\mathrm{eff}}$ saturates at $6.6$ vs. theoretical $1/\rho=6.7$ (synthetic); at $2.2$ vs. $2.0$ (LLM). Uniform inspection (Thm. 6.7) $H=20$, $\eta=0.9$, $m=3$ checkpoints Uniform spacing achieves 40–47% lower worst-case error than front-loaded, back-loaded, or random schedules. Critical horizon (Cor. 3.9) Synthetic, $\eta\in\{0.7,0.8\}$, n=1,000n=1{,}000 Attribution accuracy drops to chance ($50\%$) at or before the predicted $H_{\mathrm{crit}}$. Objective mismatch (Def. 5.1) 50 LLM reasoning chains on GSM8K 16% of chains are “mostly correct” (additive reward $\geq 0.80$) yet produce wrong answers.
8 Conclusion
This paper establishes information-theoretic limits on credit assignment in deep sequential processes, revealing a sharp sample-complexity threshold that constrains both inspection system design and reasoning model training.
The core insight.
The Signal Decay Bound (Theorem 3.8) shows that information about step $t$ decays as $\eta^{H-t}$, defining a critical horizon $H_{\mathrm{crit}}\approx\ln(n\Delta^{2})/\ln(1/\eta)$ (up to the $\epsilon$-correction in Corollary 3.9). Steps within $H_{\mathrm{crit}}$ of the outcome can be evaluated from endpoint data; steps beyond require exponentially many samples in the distance HtH-t.
Why this matters for AI reasoning.
The framework explains why outcome-trained value networks collapse to near-constant predictions on long reasoning chains (Kazemnejad et al. 2024): this is statistically optimal behavior when outcome labels carry negligible information about early steps. Token-level dynamics are deterministic, but credit assignment targets semantic content, and the many-to-one abstraction from tokens to meaning is where contraction arises (see Section 18 in the Electronic Companion for a detailed treatment). Process reward models work by providing signal beyond the critical horizon; width-based methods like GRPO (DeepSeek-AI 2024) face diminishing returns because correlation caps effective width at $1/\rho$ (Theorem 4.3).
Beyond AI: unifying inspection design.
The same mathematics applies across operations settings (manufacturing, service operations, supply chains) wherever sequential processes are evaluated at endpoints. The critical horizon formula provides a direct diagnostic for whether intermediate inspection is necessary, and the optimal scheduling results (Section 6) show that uniform spacing is minimax-optimal under homogeneous contraction, with a greedy algorithm handling heterogeneous contraction. Achieving polynomial sample complexity requires inspection density scaling as $\Omega(H/\log H)$ (Corollary 6.14). A separate failure mode arises from objective structure: standard additive reward aggregation optimizes expected correct steps rather than end-to-end validity (Section 5), and the curriculum of Proposition 5.3 bridges this gap by interpolating between learnable and correct objectives as policy quality improves.
Limitations and future directions.
Key extensions include data-driven estimation of $\eta$ and $\Delta^{2}$, learning the abstraction $\phi$ from observations, and moving from binary to graded validity.
By establishing what endpoint data can and cannot reveal, this work provides a common analytical foundation for inspection design in operations and supervision design in AI.
9 Code and Data Disclosure
The code and data to support the numerical experiments in this paper will be made publicly available upon acceptance.